### v1

# arXiv:2603.02592v1[math.AG] 03 Mar 2026

Abstract
Let $C$ be a smooth complex projective curve of genus $g>1$ and $A=J(C)$ its Jacobian with principal polarization $\Theta$. Starting from a semistable vector bundle $V$ on $C$ with $\mu(V)>2g-2$, we consider the Fourier–Mukai transform $E=\Phi_{\mathcal{P}}(a_{*}V)$. We prove that $E(\Theta)$ satisfies the $\mathrm{IT}_{0}$ property. As a consequence, for every $m\geq g+1$, the polarized Jacobian $(A,m\Theta)$ admits Ulrich bundles constructed functorially from $V$. Further we analyze stability and Chern classes of the resulting bundles. We show that the construction induces a natural morphism from the generically finite cover of moduli space of stable bundles on $C$ to the moduli space of stable bundles on $A$, producing positive-dimensional families of stable Ulrich bundles.
1 Introduction
Ulrich bundles occupy a distinguished position in the geometry of projective embeddings. They may be viewed as extremal objects with respect to cohomological vanishing and linear resolutions. They play an important role in the study of syzygies, Chow forms, and derived categories [2, 6]. A central problem is to determine which polarized varieties admit Ulrich bundles and to understand their geometric structure when they exist.
The existence question has been resolved in several important cases, including hypersurfaces, complete intersections, and certain classes of surfaces. In particular, Beauville proved the existence of Ulrich bundles on abelian surfaces [1].In higher dimensions, however, systematic constructions remain comparatively sparse, even for classical varieties such as Jacobians.
Beauville’s construction exploits special features of abelian surfaces and does not extend formally to higher-dimensional Jacobians. The present work provides a uniform construction in all dimensions, governed by slope positivity on the curve and Fourier–Mukai duality. In particular, the resulting Ulrich bundles arise functorially from stable bundles on $C$. Our construction produces positive-dimensional families in moduli space, linking the geometry of vector bundles on $C$ to Ulrich theory on $J(C)$. To the best of our knowledge, this provides the first general construction of Ulrich bundles on Jacobians in arbitrary dimension, together with a geometric description of the induced families in moduli.
Let $C$ be a smooth complex projective curve of genus $g>1$, and let
$A=J(C)$
be its Jacobian endowed with its principal polarization $\Theta$. Although the geometry of $A$ is well understood, a uniform construction of Ulrich bundles on the polarized varieties $(A,m\Theta)$ has not previously been carried out in this generality.
The purpose of this paper is to provide such a construction. Our starting point is a semistable vector bundle $V$ on $C$ satisfying the slope condition
$\mu(V)>2g-2.$
Let
$a:C\hookrightarrow A$
be the Abel–Jacobi embedding and let $\mathcal{P}$ denote the normalized Poincaré bundle on $A\times\widehat{A}$. We consider the Fourier–Mukai transform (see subsection 2.3)
$E:=\Phi_{\mathcal{P}}(a_{*}V).$
The key observation is that positivity of slope on the curve translates into strong cohomological control on the Jacobian. More precisely, the inequality $\mu(V)>2g-2$ implies uniform vanishing of $H^{1}(C,V\otimes M)$ for all $M\in Pic^{0}(C)$. Via derived base change, this yields the $\mathrm{WIT}_{0}$ property for $a_{*}V$ and shows that $E$ is locally free with explicitly computable rank. We then prove that the twist $E(\Theta)$ satisfies the stronger $\mathrm{IT}_{0}$ property by reducing cohomology on the Jacobian directly to cohomology on the curve.
The $\mathrm{IT}_{0}$ property implies continuous global generation in the sense of Pareschi–Popa [7]. A kernel construction then produces Ulrich bundles on $(A,m\Theta)$ for all sufficiently positive polarizations.
Theorem A (See Theorem 5.2). Let $C$ be a smooth projective curve of genus $g>1$ and let $V$ be a semistable vector bundle on $C$ with $\mu(V)>2g-2$. Then for every integer $m\geq g+1$, the polarized Jacobian $(J(C),m\Theta)$ admits an Ulrich bundle constructed functorially from $V$.
The rank of the resulting Ulrich bundle is determined by $h^{0}(C,V)$. Specializing to line bundles of degree >2g2>2g-2 yields an upper bound for the Ulrich complexity of $(J(C),m\Theta)$.
Beyond existence, we analyze structural properties of the constructed bundles. Under natural stability assumptions on $V$, we prove that the resulting Ulrich bundles are slope-stable for large polarizations. Using Grothendieck–Riemann–Roch, we compute their Chern classes and show that all bundles arising in the construction have first Chern class proportional to the principal polarization $\Theta$.
The construction is compatible with families and has geometric consequences at the level of moduli space. Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$. The above construction induces a natural morphism from a generically finite cover of $M_{C}$ to the moduli space of stable bundles on $A$ with the corresponding numerical invariants.
This compatibility with families has strong geometric consequences. In particular, the Ulrich locus inside the moduli space of bundles on A is highly nontrivial:
Theorem B (See Theorem 10.2)
Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$. For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem 5.2 Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
$\pi:T\longrightarrow M_{C},$
and a dense open subset $T^{\mathrm{st}}\subset T$ such that the Ulrich construction is compatible with families and induces a morphism
$\Phi:T^{\mathrm{st}}\longrightarrow M_{A}.$
In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$.
Thus the existence of Ulrich bundles on $(A,m\Theta)$ is governed by slope positivity on the curve and gives rise to a geometric link between vector bundles on $C$ and stable bundles on its Jacobian.
Layout of the paper. Section 2 reviews background on slope stability and Fourier–Mukai transforms. Section 3 establishes the $\mathrm{WIT}_{0}$ property and computes the rank of the transform. Section 4 proves that $E(\Theta)$ satisfies $\mathrm{IT}_{0}$. Section 5 constructs Ulrich bundles and proves Theorem A. Section 6 discusses the numerical bound $m\geq g+1$. Section 7 studies slope stability. Section 8 computes Chern classes. Section 9 specializes to line bundles and derives an upper bound for Ulrich complexity. Section 10 constructs the moduli morphism and proves Theorem B.
2 Preliminaries
2.1 Ulrich bundles and Ulrich complexity
Let $(X,H)$ be a polarized smooth projective variety of dimension $n$, where $H$ is an ample divisor on $X$. A vector bundle $E$ on $X$ is called an Ulrich bundle with respect to $H$ if
$H^{i}(X,E(-jH))=0\qquad\text{for all }i\text{ and }1\leq j\leq n.$
In this paper we will consider Ulrich bundles on polarized Jacobians
$(A,m\Theta),$
where $A=J(C)$ is the Jacobian of a smooth complex projective curve $C$ of genus $g>1$ and $\Theta$ denotes the principal polarization.
Ulrich complexity. For a polarized variety $(X,H)$ admitting Ulrich bundles, we define the Ulrich complexity to be
$\mathrm{uc}(X,H):=\min\{rk(E)\;:\;E\text{ is an Ulrich bundle on }(X,H)\}.$
In general, determining $\mathrm{uc}(X,H)$ is a subtle problem. In the present work, we construct explicit Ulrich bundles on $(J(C),m\Theta)$ and derive upper bounds for $\mathrm{uc}(J(C),m\Theta)$ in terms of the genus $g$ and the degree of the initial bundle on $C$.
2.2 The Abel–Jacobi embedding
Fix a base point $x_{0}\in C$. The Abel–Jacobi morphism
$a:C\longrightarrow A,\qquad x\longmapsto\mathcal{O}_{C}(x-x_{0}),$
is a closed immersion. We identify $C$ with its image in $A$.
A classical result (see [4, Ch. VII]) states that
$a^{*}\Theta\sim\text{a divisor of degree }g\text{ on }C.$
In particular,
$\deg(a^{*}\mathcal{O}_{A}(\Theta))=g.$
2.3 Fourier–Mukai transform on the Jacobian
Let $\widehat{A}:=Pic^{0}(A)$ be the dual abelian variety. Since $A$ is principally polarized, we identify
$A\cong\widehat{A}$
via $\Theta$.
Let
$\mathcal{P}$
be the normalized Poincaré bundle on $A\times\widehat{A}$, characterized by the properties:
  1. 1.
    $\mathcal{P}|_{A\times\{0\}}\cong\mathcal{O}_{A}$,
  2. 2.
    $\mathcal{P}|_{\{0\}\times\widehat{A}}\cong\mathcal{O}_{\widehat{A}}$.
Let $p_{1}$ and $p_{2}$ denote the projections from $A\times\widehat{A}$ to the first and second factor. The Fourier–Mukai functor
$Db(A^)\Phi_{\mathcal{P}}$$\longrightarrow D^{b}(\widehat{A})$
is defined by
$\Phi_{\mathcal{P}}(\mathcal{F}):=Rp_{2*}(p_{1}^{*}\mathcal{F}\otimes\mathcal{P}).$
2.4 WIT and IT conditions
Let $\mathcal{F}$ be a coherent sheaf on $A$.
Definition 2.1.
We say that $\mathcal{F}$ satisfies WITi (Weak Index Theorem of index $i$) if
$R^{j}\Phi_{\mathcal{P}}(\mathcal{F})=0\quad\text{for all }j\neq i.$
If $\mathcal{F}$ satisfies WITi, we denote
$\widehat{\mathcal{F}}:=R^{i}\Phi_{\mathcal{P}}(\mathcal{F}).$
Definition 2.2.
A coherent sheaf $\mathcal{E}$ on $A$ satisfies IT0 (Index Theorem with index 0) if
$H^{i}(A,\mathcal{E}\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in Pic^{0}(A).$
IT0 is stronger than WIT0. If $\mathcal{E}$ is locally free and satisfies IT0, then $\mathcal{E}$ is continuously globally generated in the sense of [7].
2.5 Base change for Fourier–Mukai
Let $\mathcal{F}$ be a coherent sheaf on $A$. For $\alpha\in Pic^{0}(A)$, derived base change gives
$\Phi_{\mathcal{P}}(\mathcal{F})\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(A,\mathcal{F}\otimes\mathcal{P}_{\alpha}).$
This identity will be used repeatedly to compute fibers of Fourier–Mukai transforms.
The preceding notions provide the foundational framework for the construction carried out in the subsequent sections.
2.6 Slope Stability
Let $(X,H)$ be a smooth projective variety of dimension $n$ with polarization $H$. For a torsion-free coherent sheaf $\mathcal{E}$ on $X$, the slope with respect to $H$ is
$\mu_{H}(\mathcal{E}):=\frac{c_{1}(\mathcal{E})\cdot H^{n-1}}{\operatorname{rk}(\mathcal{E})}.$
Definition 2.3.
  1. 1.
    The sheaf $\mathcal{E}$ is called $\mu_{H}$-stable if for every proper nonzero subsheaf $\mathcal{F}\subset\mathcal{E}$ with $0<\operatorname{rk}(\mathcal{F})<\operatorname{rk}(\mathcal{E})$, one has
    $\mu_{H}(\mathcal{F})<\mu_{H}(\mathcal{E}).$
  2. 2.
    The sheaf $\mathcal{E}$ is called $\mu_{H}$-semistable if the above inequality is replaced by
    $\mu_{H}(\mathcal{F})\leq\mu_{H}(\mathcal{E}).$
  3. 3.
    The sheaf $\mathcal{E}$ is called $\mu_{H}$-polystable if it is a direct sum of $\mu_{H}$-stable sheaves all having the same slope.
(See [12, Ch. 1] )
In particular, a direct sum of $\mu_{H}$-stable sheaves of identical slope is $\mu_{H}$-polystable. For example, on an abelian variety, a direct sum of algebraically trivial line bundles is $\mu_{H}$-polystable of slope zero.
Curves. If $C$ is a smooth projective curve, then
$\mu(V)=\frac{\deg(V)}{\operatorname{rk}(V)}$
for any vector bundle $V$ on $C$, and semistability is independent of a choice of polarization. Semistability is preserved under tensoring with line bundles of degree zero.
2.7 Semihomogeneous bundles on abelian varieties
Let $A$ be an abelian variety and denote by $t_{x}:A\to A$ the translation by a point $x\in A$.
Definition 2.4.
A vector bundle $\mathcal{E}$ on $A$ is called semihomogeneous if for every point $x\in A$ there exists a line bundle $\alpha_{x}\in Pic^{0}(A)$ such that
$t_{x}^{*}\mathcal{E}\cong\mathcal{E}\otimes\alpha_{x}.$
Semihomogeneous bundles were introduced and studied by Mukai [15]. They enjoy the following properties:
  • A semihomogeneous bundle is simple if and only if it is stable with respect to any polarization [15, Prop. 6.16].
  • If $\mathcal{E}$ is semihomogeneous, then its Chern classes are proportional to powers of $c_{1}(\mathcal{E})$ in the Néron–Severi group [15, Thm. 5.8].
3 Cohomological Vanishing and the WIT0 Property
In this section we analyze the Fourier–Mukai transform of the pushforward $a_{*}V$ of a semistable vector bundle $V$ on $C$. The goal is to establish uniform cohomological vanishing on $C$, derive the WIT0 property for $a_{*}V$, and prove that its Fourier–Mukai transform is locally free with explicitly computable rank.
3.1 Slope positivity and uniform $H^{1}$-vanishing
Let $V$ be a vector bundle on $C$. Recall that $V$ is semistable if
$\mu(W)\leq\mu(V)\quad\text{for every proper subbundle }W\subset V.$
We assume throughout this section that
$\mu(V)>2g-2.$
Lemma 3.1 (Uniform $H^{1}$-vanishing).
If $V$ is semistable and $\mu(V)>2g-2$, then
$H^{1}(C,V\otimes M)=0\quad\text{for all }M\in Pic^{0}(C).$
Proof.
For $M\in Pic^{0}(C)$, the bundle $V\otimes M$ remains semistable and satisfies
$\mu(V\otimes M)=\mu(V).$
By Serre duality,
$H^{1}(C,V\otimes M)\cong\operatorname{Hom}(V\otimes M,\omega_{C})^{\vee}.$
Since $\omega_{C}$ is a line bundle of slope 2g22g-2, and $\mu(V)>2g-2$, semistability implies
$\operatorname{Hom}(V\otimes M,\omega_{C})=0.$
The claim follows. ∎
Remark 3.2.
The slope inequality $\mu(V)>2g-2$ is sharp for this argument. It ensures that $V$ admits no nonzero morphism to $\omega_{C}$, which is precisely the obstruction to the vanishing of $H^{1}$.
3.2 Derived fibers and WIT0
Let
$E:=\Phi_{\mathcal{P}}(a_{*}V).$
We now compute the derived fiber of $E$ at a point $\alpha\in Pic^{0}(A)$.
By derived base change for the Fourier–Mukai transform,
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(A,a_{*}V\otimes\mathcal{P}_{\alpha}).$
Using the projection formula and $R\Gamma(A,a_{*}(-))\cong R\Gamma(C,-)$, we obtain
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(C,V\otimes a^{*}\mathcal{P}_{\alpha}).$
Since $a^{*}\mathcal{P}_{\alpha}\in Pic^{0}(C)$, Lemma 3.1 implies that
$H^{1}(C,V\otimes a^{*}\mathcal{P}_{\alpha})=0\quad\forall\alpha.$
Thus the derived fiber is concentrated in degree 0.
Proposition 3.3 (WIT0 and local freeness).
Assume $\mu(V)>2g-2$. Then:
  1. 1.
    $a_{*}V$ satisfies WIT0 with respect to $\Phi_{\mathcal{P}}$;
  2. 2.
    $E$ is a locally free sheaf on $A$;
  3. 3.
    $\operatorname{rk}(E)=h^{0}(C,V)$.
Proof.
Since for every $\alpha$ the derived fiber $E\otimes^{\mathbf{L}}k(\alpha)$ is concentrated in degree 0, we conclude that
$R^{i}\Phi_{\mathcal{P}}(a_{*}V)=0\quad\text{for all }i>0.$
Thus $a_{*}V$ satisfies WIT0.
Moreover,
$E\otimes k(\alpha)\simeq H^{0}(C,V\otimes a^{*}\mathcal{P}_{\alpha}),$
and the dimension of this space is constant in $\alpha$ by semicontinuity. Hence $E$ is locally free.
Finally,
$\operatorname{rk}(E)=h^{0}(C,V).$
3.3 Rank formula
Applying Riemann–Roch on $C$ and using Lemma 3.1, we obtain:
Corollary 3.4 (Rank formula).
If $\mu(V)>2g-2$, then
$\operatorname{rk}(E)=\deg(V)+\operatorname{rk}(V)(1-g).$
Proof.
Since $H^{1}(C,V)=0$, Riemann–Roch gives
$h^{0}(C,V)=\deg(V)+\operatorname{rk}(V)(1-g).$
The result follows from Proposition $3.$ ∎
The results of this section show that positivity of slope on the curve translates into strong structural properties of the Fourier–Mukai transform on the Jacobian. In particular, $E$ is a well-behaved vector bundle whose rank and fibers admit explicit description.
The next section upgrades this WIT0 property to an IT0 statement by a direct cohomological reduction to the curve.
4 From WIT0 to IT0
In Section 3 we proved that
$E:=\Phi_{\mathcal{P}}(a_{*}V)$
is locally free and that $a_{*}V$ satisfies WIT0. In order to construct Ulrich bundles, we now establish directly that a suitable twist of $E$ satisfies IT0.
We prove the required cohomological vanishing by an explicit reduction to the curve.
4.1 Reduction to cohomology on the curve
Let $\alpha\in\mathrm{Pic}^{0}(A)$. By definition of the Fourier–Mukai transform and derived base change (see Mukai [3, 5]), we have
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(A,a_{*}V\otimes\mathcal{P}_{\alpha}).$
Using the projection formula and the identity $R\Gamma(A,a_{*}(-))\simeq R\Gamma(C,-)$, we obtain
$E\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(C,V\otimes a^{*}\mathcal{P}_{\alpha}).$
Twisting by $\Theta$ and taking cohomology yields
$H^{i}(A,E(\Theta)\otimes\alpha)\simeq H^{i}(C,V\otimes a^{*}\mathcal{O}_{A}(\Theta)\otimes a^{*}\mathcal{P}_{\alpha}).$
Since $a^{*}\mathcal{P}_{\alpha}\in\mathrm{Pic}^{0}(C)$, this reduces the IT0 condition for $E(\Theta)$ to uniform vanishing statements on the curve.
4.2 Slope computation
Recall that
$\deg(a^{*}\mathcal{O}_{A}(\Theta))=g.$
Hence
$\mu\!\left(V\otimes a^{*}\mathcal{O}_{A}(\Theta)\right)=\mu(V)+g.$
Under the standing assumption
$\mu(V)>2g-2,$
we obtain
$\mu(V)+g>3g-2>2g-2.$
Therefore the bundle
$V\otimes a^{*}\mathcal{O}_{A}(\Theta)$
is semistable of slope strictly greater than 2g22g-2.
4.3 Vanishing of $H^{1}$
By Lemma 3.1 applied to
$V\otimes a^{*}\mathcal{O}_{A}(\Theta),$
we obtain
$H^{1}\!\left(C,V\otimes a^{*}\mathcal{O}_{A}(\Theta)\otimes M\right)=0\quad\forall\,M\in\mathrm{Pic}^{0}(C).$
In particular,
$H^{1}(A,E(\Theta)\otimes\alpha)=0\quad\forall\alpha\in\mathrm{Pic}^{0}(A).$
4.4 Vanishing of higher cohomology
Since $C$ has dimension $1$, we have
$H^{i}(C,-)=0\quad\text{for }i\geq 2.$
Therefore
$H^{i}(A,E(\Theta)\otimes\alpha)=0\quad\forall i\geq 2.$
4.5 Conclusion
Combining the above vanishing results, we obtain:
Theorem 4.1 (IT0 property of $E(\Theta)$).
Let $V$ be a semistable vector bundle on $C$ with $\mu(V)>2g-2$. Then
$E(\Theta)$
satisfies IT0, i.e.
$H^{i}(A,E(\Theta)\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in\mathrm{Pic}^{0}(A).$
Proof.
The vanishing for $i=1$ follows from the slope computation and Lemma 3.1. The vanishing for $i\geq 2$ follows from the fact that cohomology on the curve vanishes in degrees $\geq 2$. ∎
5 Construction of Ulrich Bundles
In this section we complete the proof of Theorem A by constructing Ulrich bundles on $(A,m\Theta)$ from the IT0 property established in Section 4.
Throughout this section we assume that $V$ is semistable with $\mu(V)>2g-2$, and we set
$E:=\Phi_{\mathcal{P}}(a_{*}V).$
By Theorem 4.1,
$F:=E(\Theta)$
satisfies IT0.
5.1 Positivity and higher cohomology vanishing
We first record a simple consequence of the IT0 property.
Lemma 5.1.
Let $F=E(\Theta)$ be as above. If $F$ satisfies IT0, then for every integer $k\geq 1$,
$H^{i}(A,F(k\Theta)\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in\mathrm{Pic}^{0}(A).$
In particular,
$H^{i}(A,E((k+1)\Theta))=0\quad\forall i>0,\ k\geq 1.$
Proof.
Since $F$ satisfies IT0, we have
$H^{i}(A,F\otimes\beta)=0\quad\forall i>0,\ \forall\beta\in\mathrm{Pic}^{0}(A).$
For $k\geq 1$, the line bundle $\mathcal{O}_{A}(k\Theta)$ is ample. On an abelian variety, twisting an IT0 sheaf by an ample line bundle preserves the vanishing of higher cohomology (see [7, Thm. 2.3]).
Hence
$H^{i}(A,F(k\Theta)\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in\mathrm{Pic}^{0}(A).$
Taking $\alpha=\mathcal{O}_{A}$ gives the second statement. ∎
5.2 Continuous global generation
A fundamental consequence of the IT0 property is continuous global generation.
By [7, 8], a locally free sheaf satisfying IT0 is continuously globally generated. Hence there exist finitely many
$\alpha_{1},\dots,\alpha_{N}\in\mathrm{Pic}^{0}(A)$
such that the evaluation morphism
$\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}\longrightarrow F$
is surjective.
Let
$0\longrightarrow M\longrightarrow G\longrightarrow F\longrightarrow 0$
be the resulting exact sequence, where
$G:=\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}.$
5.3 Twisting and Ulrich vanishing
Let $H=m\Theta$ with $m\geq g+1$, and define
$U:=M(H).$
We verify that $U$ is Ulrich with respect to $H$.
The vector bundle $U$ on $(A,H)$ is Ulrich if
$H^{i}(A,U(-pH))=0\quad\forall i,\ 1\leq p\leq g.$
For $1\leq p\leq g$, twist the above evaluation sequence by (HpH)(H-pH):
$0\to M(H-pH)\to G(H-pH)\to F(H-pH)\to 0.$
Since $U=M(H)$, we have
U(pH)=M(HpH).U(-pH)=M(H-pH).
We analyze the two terms.
Vanishing for G(HpH)G(H-pH).
Each summand of $G$ is of the form $\alpha_{k}^{-1}$. Hence
$G(H-pH)=\bigoplus_{k}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}((m-p)\Theta).$
If $1\leq p\leq g$ and $m\geq g+1$, then $(m-p)\Theta$ is ample. On an abelian variety, an ample line bundle satisfies
$H^{i}(A,\mathcal{O}_{A}((m-p)\Theta)\otimes\alpha)=0\quad\forall i>0.$
Thus
$H^{i}(A,G(H-pH))=0\quad\forall i>0.$
Vanishing for F(HpH)F(H-pH).
We compute
$F(H-pH)=E(\Theta+(m-p)\Theta)=E((m-p+1)\Theta).$
Since $m\geq g+1$ and $1\leq p\leq g$, we have
$m-p+1\geq 2.$
Thus F(HpH)F(H-pH) is a positive twist of $F$. By Lemma 5.1,
$H^{i}(A,F(H-pH))=0\quad\forall i>0.$
5.4 Conclusion
The long exact sequence in cohomology associated to the twisted evaluation sequence yields
$H^{i}(A,U(-pH))=0\quad\forall i,\ 1\leq p\leq g.$
Hence $U$ is an Ulrich bundle on $(A,m\Theta)$.
Hence we conclude the main theorem
Theorem 5.2.
Let $C$ be a smooth projective curve of genus $g>1$ and let $V$ be a semistable vector bundle on $C$ with $\mu(V)>2g-2$. Then for every integer $m\geq g+1$, the polarized Jacobian $(J(C),m\Theta)$ admits an Ulrich bundle constructed functorially from $V$.
Remark 5.3.
We do not claim optimality of the bound $m\geq g+1$. The construction is functorial in $V$, and the rank of the resulting Ulrich bundle depends explicitly on $h^{0}(C,V)$.
6 On the bound $m\geq g+1$
In Theorem 5.2 we obtained Ulrich bundles on $(A,m\Theta)$ for all integers $m\geq g+1$. We briefly explain why this numerical condition naturally arises from the construction.
Recall that $U=M(m\Theta)$, where
$0\to M\to G\to F\to 0,\qquad F=E(\Theta),$
and $F$ satisfies $\mathrm{IT}_{0}$. To verify the Ulrich condition, one must prove that
$H^{i}(A,U(-pH))=0\quad\text{for all }i\text{ and }1\leq p\leq g,$
where $H=m\Theta$.
Since
$U(-pH)=M((m-p)\Theta),$
we twist the evaluation sequence by $(m-p)\Theta$. The argument in Section 5 uses two facts:
  1. (i)
    $(m-p)\Theta$ must be ample to ensure vanishing for the bundle $G$;
  2. (ii)
    the twist $E((m-p+1)\Theta)$ must remain a positive twist of $E(\Theta)$ so that Lemma 5.1 applies.
Because $\dim A=g$, the most restrictive case is $p=g$. Thus we must have
$m-g+1\geq 2,$
which is equivalent to
$m\geq g+1.$
If $m=g$, then the extremal term involves $E(\Theta)$ itself, and the argument does not guarantee the required vanishing. If $m<g$, some twists $(m-p)\Theta$ cease to be ample, and the cohomological vanishing used in Section 5 breaks down.
Therefore the condition $m\geq g+1$ is intrinsic to the present method. We do not claim optimality of this bound; the existence of Ulrich bundles for smaller values of $m$ remains open.
7 Stability of the Ulrich Bundle
In this section we study slope stability of the Ulrich bundles constructed in Section 5.
Throughout, let $C$ be a smooth projective curve of genus $g>1$, let $A=J(C)$ be its Jacobian with principal polarization $\Theta$, and let $V$ be a stable vector bundle on $C$ with $\mu(V)>2g-2$. Set
$E:=\Phi_{\mathcal{P}}(a_{*}V),\qquad F:=E(\Theta).$
By Theorem 4.1, $F$ satisfies IT0.
Stability of $F$
We first see that Fourier–Mukai transforms of stable bundles supported on a curve often yield simple semihomogeneous bundles on the Jacobian.
Proposition 7.1.
Assume that $V$ is stable on $C$. Then $F=E(\Theta)$ is a simple semihomogeneous vector bundle on $A$. In particular, $F$ is $\mu_{\Theta}$-stable.
Proof.
We first prove simplicity. Since $a:C\hookrightarrow A$ is a closed immersion, the pushforward functor $a_{*}$ is exact and fully faithful on coherent sheaves supported on $C$; in particular,
$Hom_{A}(a_{*}V,a_{*}V)\;\cong\;Hom_{C}(V,V).$
Because $V$ is stable on a curve, $End_{C}(V)=\mathbb{C}$ (see [12, Ch. 1]), hence $Hom_{A}(a_{*}V,a_{*}V)=\mathbb{C}$.
Now we apply the Fourier–Mukai equivalence $\Phi_{\mathcal{P}}:D^{b}(A)\to D^{b}(A)$. Since $\Phi_{\mathcal{P}}$ is an equivalence, it preserves $Hom$ in the derived category:
$Hom_{D^{b}(A)}(a_{*}V,a_{*}V)\;\cong\;Hom_{D^{b}(A)}(\Phi_{\mathcal{P}}(a_{*}V),\Phi_{\mathcal{P}}(a_{*}V)).$
By Proposition 3.3 in this paper, $E=\Phi_{\mathcal{P}}(a_{*}V)$ is a vector bundle, hence $Hom_{D^{b}(A)}(E,E)=Hom_{A}(E,E)$. Combining with the previous paragraph gives $End_{A}(E)=\mathbb{C}$, i.e. $E$ is simple; therefore $F=E(\Theta)$ is also simple.
Next we explain semihomogeneity. The sheaf $E=\Phi_{\mathcal{P}}(a_{*}V)$ is a Picard sheaf in the sense of Mukai (Fourier–Mukai transform of a sheaf pushed forward from a curve inside its Jacobian). Mukai proves that Picard sheaves on abelian varieties are semihomogeneous; see [3, §3–§4] (and also the exposition [14, Ch. 6]). Thus $E$ is semihomogeneous, and twisting by a line bundle preserves semihomogeneity, so $F=E(\Theta)$ is semihomogeneous as well.
Finally, by Mukai’s stability criterion, a simple semihomogeneous vector bundle on an abelian variety is $\mu_{H}$-stable with respect to every polarization $H$; see [5, Prop. 6.16] (or [14, Prop. 6.16]). Taking $H=\Theta$ yields that $F$ is $\mu_{\Theta}$-stable. ∎
Slope of the evaluation bundle
We recall from Section 5 that, since $F$ satisfies IT0, it is continuously globally generated. Thus there exist finitely many $\alpha_{1},\dots,\alpha_{N}\in\mathrm{Pic}^{0}(A)$ and a surjection
$\varphi:\;G:=\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}\longrightarrow F.$
Let
$0\longrightarrow M\longrightarrow G\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}F\longrightarrow 0$
be the corresponding exact sequence.
Lemma 7.2.
$G$ is $\mu_{\Theta}$-polystable of slope 0.
Proof.
Each summand $\alpha_{k}^{-1}$ lies in $\mathrm{Pic}^{0}(A)$ and thus has trivial first Chern class. Hence $c_{1}(\alpha_{k}^{-1})=0$, so $\mu_{\Theta}(\alpha_{k}^{-1})=0$. Being a direct sum of degree-zero line bundles, $G$ is polystable of slope 0. ∎
Since $F$ is $\mu_{\Theta}$-stable and $c_{1}(F)$ is proportional to $\Theta$, we have
$\mu_{\Theta}(F)>0.$
Stability of the kernel
We now prove stability of the kernel for a general choice of evaluation map.
Proposition 7.3.
Assume $F$ is $\mu_{\Theta}$-stable. For a general surjection $\varphi:G\twoheadrightarrow F$ as above, the kernel $M=\ker(\varphi)$ is $\mu_{\Theta}$-stable.
Proof.
Suppose that $M$ is not stable. Then there exists a proper saturated subsheaf $K\subset M$ with
$\mu_{\Theta}(K)\geq\mu_{\Theta}(M).$
Composing $K\hookrightarrow M\hookrightarrow G$, we obtain a nonzero subsheaf of $G$. Since $G$ is polystable of slope 0, any subsheaf has slope $\leq 0$.
On the other hand, from the exact sequence we compute
$\mu_{\Theta}(M)=\frac{rk(G)\mu_{\Theta}(G)-rk(F)\mu_{\Theta}(F)}{rk(M)}=-\frac{rk(F)}{rk(M)}\,\mu_{\Theta}(F)<0.$
Thus any destabilizing $K$ would satisfy
$\mu_{\Theta}(K)\geq\mu_{\Theta}(M)>-\mu_{\Theta}(F).$
A standard parameter-count argument in $\operatorname{Hom}(G,F)$ shows that the existence of such a $K$ forces $\varphi$ to lie in a proper closed subset of $\operatorname{Hom}(G,F)$. Hence for a general choice of $\varphi$, no such destabilizing $K$ exists. Therefore $M$ is $\mu_{\Theta}$-stable. ∎
Stability of the Ulrich bundle
Finally, let $H=m\Theta$ with $m\geq g+1$ and define
$U:=M(H).$
Theorem 7.4.
Under the above assumptions, for a general choice of evaluation map and for all $m\gg 0$, the Ulrich bundle $U$ on $(A,m\Theta)$ is $\mu_{m\Theta}$-stable.
Proof.
Twisting by a line bundle preserves slope stability. Hence $U$ is $\mu_{m\Theta}$-stable if and only if $M$ is $\mu_{m\Theta}$-stable.
Stability is an open condition with respect to the polarization ray $\mathbb{R}_{>0}\Theta$. Since $M$ is $\mu_{\Theta}$-stable for general $\varphi$, it remains stable for $m\gg 0$. ∎
8 Chern Class
In this section we compute the Chern character of the Fourier–Mukai bundle
$E=\Phi_{\mathcal{P}}(a_{*}V)$
and determine the first Chern classes of the bundles appearing in our Ulrich construction. These computations clarify the numerical structure of the construction and will be useful for slope considerations.
Geometric setup
Let $C$ be a smooth projective curve of genus $g>1$, let $A=J(C)$ be its Jacobian with principal polarization $\Theta$, and let $a:C\hookrightarrow A$ denote the Abel–Jacobi embedding.
Let $V$ be a vector bundle on $C$ of rank $r$ and degree $d$. We assume throughout that $\mu(V)>2g-2$, so that the Fourier–Mukai transform
$E=\Phi_{\mathcal{P}}(a_{*}V)$
is locally free (Section 3).
Recall that
$E=Rp_{2*}\!\left(p_{1}^{*}a_{*}V\otimes\mathcal{P}\right),$
where $\mathcal{P}$ is the normalized Poincaré bundle on $A\times A$ and $p_{1},p_{2}$ are the projections. Since $a_{*}V$ is supported on the embedded curve $C\subset A$, this expression reduces to a pushforward along
$p_{2}:C\times A\longrightarrow A.$
Applying Grothendieck–Riemann–Roch
We apply Grothendieck–Riemann–Roch to the morphism $p_{2}$. It gives
$\mathrm{ch}(E)=p_{2*}\left(\mathrm{ch}(p_{1}^{*}V)\cdot\mathrm{ch}(\mathcal{P}|_{C\times A})\cdot\mathrm{td}(C)\right).$
Since $C$ has dimension one, the Chern character and Todd class simplify considerably:
$\mathrm{ch}(V)=r+d\,[\mathrm{pt}],\qquad\mathrm{td}(C)=1+(1-g)\,[\mathrm{pt}].$
Thus
$\mathrm{ch}(V)\cdot\mathrm{td}(C)=r+(d+r(1-g))[\mathrm{pt}].$
The Poincaré bundle satisfies
$\mathrm{ch}(\mathcal{P})=1+c_{1}(\mathcal{P})+\tfrac{1}{2}c_{1}(\mathcal{P})^{2}+\cdots,$
where
$c_{1}(\mathcal{P})=\sum_{i}\alpha_{i}\otimes\beta_{i},$
with $\{\alpha_{i}\}$ a basis of $H^{1}(C)$ and $\{\beta_{i}\}$ the dual basis of $H^{1}(A)$ (see [14, Ch. 6]).
Only terms of total codimension two on $C\times A$ contribute to the first Chern class of $E$ after pushing forward.
Rank of $E$
The rank of $E$ comes from the degree-one term on $C$:
rk(E)=d+r(1g).rk(E)=d+r(1-g).
Since our slope assumption implies $H^{1}(C,V)=0$, Riemann–Roch gives
$h^{0}(C,V)=d+r(1-g),$
so we recover the rank formula obtained earlier in Proposition 3.3:
$rk(E)=h^{0}(C,V).$
First Chern class of $E$
To compute $c_{1}(E)$, we isolate the codimension-two contribution coming from
$r\cdot c_{1}(\mathcal{P})^{2}.$
A classical computation (see [14, Ch. 6] and [5, Prop. 6.16]) shows that
$p_{2*}\!\left(c_{1}(\mathcal{P})^{2}\right)=2\,\Theta.$
Consequently,
$c_{1}(E)=r\,\Theta.$
Thus the determinant of $E$ is a positive multiple of the principal polarization.
First Chern class of $F=E(\Theta)$
Since twisting by $\Theta$ adds $rk(E)\Theta$ to the first Chern class, we obtain
$c_{1}(F)=c_{1}(E)+rk(E)\,\Theta=\big(r+rk(E)\big)\Theta.$
Chern classes of the kernel and the Ulrich bundle
Recall the exact sequence constructed in Section 5:
$0\longrightarrow M\longrightarrow G\longrightarrow F\longrightarrow 0,$
where $G$ is a direct sum of degree-zero line bundles. Hence
$c_{1}(G)=0,$
and therefore
$c_{1}(M)=-c_{1}(F).$
Finally, the Ulrich bundle is defined by
$U=M(m\Theta).$
Twisting adds $rk(M)m\Theta$ to the first Chern class, giving
$c_{1}(U)=c_{1}(M)+rk(M)m\Theta=\big(rk(M)m-r-rk(E)\big)\Theta.$
Remark 8.1.
In particular, all bundles arising in our construction have first Chern class proportional to the principal polarization $\Theta$. This reflects the intrinsic compatibility of the construction with the polarization of the Jacobian. It suggests that the resulting Ulrich bundles occupy a distinguished ray in the Néron–Severi space of $A$.
9 Specialization to the Line Bundle
Let $L$ be a line bundle on $C$. then $L$ be automatically semistable. Further we assume $\deg L>2g-2$. Thus we apply the above theory.
Corollary 9.1.
Let $L$ be a line bundle with $\deg L>2g-2$. Then $(J(C),m\Theta)$ carries an Ulrich bundle for all $m\geq g+1$.
9.1 An upper bound for the Ulrich complexity
In this subsection we compute rank of the Ulrich bundles produced by the construction of Section 5 in the case of line bundles.
Let $L$ be a line bundle on $C$ of degree d>2g2d>2g-2, and set
$E=\Phi_{\mathcal{P}}(a_{*}L),\qquad F=E(\Theta).$
By Theorem 4.1, $F$ satisfies IT0. In particular, $F$ is continuously globally generated (CGG) in the sense of Pareschi–Popa (see [7, Remark 2.2(b)]).
Hence there exist finitely many line bundles
$\alpha_{1},\dots,\alpha_{N}\in\operatorname{Pic}^{0}(A)$
such that the sum of the twisted evaluation maps yields a surjection
$\varphi:G:=\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}\longrightarrow F.$
Let
$0\longrightarrow M\longrightarrow G\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}F\longrightarrow 0$
be the corresponding exact sequence, and define the Ulrich bundle
$U:=M(m\Theta),\qquad m\geq g+1.$
Computation of $h^{0}(A,F\otimes\alpha)$
Since $F$ satisfies IT0, we have
$H^{i}(A,F\otimes\alpha)=0\quad\text{for all }i>0\text{ and all }\alpha\in\operatorname{Pic}^{0}(A).$
Therefore
$h^{0}(A,F\otimes\alpha)=\chi(A,F\otimes\alpha)=\chi(A,F),$
which is independent of $\alpha$.
By Riemann–Roch on the abelian variety $A$ (using $\mathrm{td}(A)=1$),
$\chi(A,F)=\frac{c_{1}(F)^{g}}{g!}.$
From Section 8 we have
$c_{1}(E)=\Theta,\qquad rk(E)=d+1-g.$
Hence
$c_{1}(F)=c_{1}(E)+rk(E)\Theta=(1+rk(E))\Theta.$
Set
n:=rk(E)=d+1g.n:=rk(E)=d+1-g.
Then
$c_{1}(F)=(1+n)\Theta.$
Since $(\Theta^{g})=g!$ for a principally polarized abelian variety, we obtain
$\chi(A,F)=\frac{(1+n)^{g}\Theta^{g}}{g!}=(1+n)^{g}.$
Thus
$h^{0}(A,F\otimes\alpha)=(1+n)^{g}\quad\text{for every }\alpha\in\operatorname{Pic}^{0}(A).$
Rank of the Ulrich bundle
By construction,
$rk(G)=\sum_{k=1}^{N}h^{0}(A,F\otimes\alpha_{k})=N(1+n)^{g}.$
Since $rk(F)=n$, it follows that
$rk(U)=rk(M)=rk(G)-rk(F)=N(1+n)^{g}-n.$
Substituting n=d+1gn=d+1-g, we obtain:
Proposition 9.2.
Let $L$ be a line bundle on $C$ of degree d>2g2d>2g-2. Then the Ulrich bundle $U$ constructed in Section 5 satisfies
$rk(U)=N(d+2-g)^{g}-(d+1-g),$
where $N$ is a positive integer depending on $F$, whose existence is guaranteed by continuous global generation. In particular,
$\mathrm{uc}(J(C),m\Theta)\leq N(d+2-g)^{g}-(d+1-g).$
Minimal degree choice
If we choose the minimal possible degree d=2g1d=2g-1, then $n=g$ and
$rk(U)=N(g+1)^{g}-g.$
Then we get an upper bound
$\mathrm{uc}(J(C),m\Theta)\leq N(g+1)^{g}-g.$
Remark 9.3.
The integer $N$ arises from the continuous global generation of $F$. While Pareschi–Popa guarantee the existence of such a finite $N$, no effective universal bound for $N$ is presently known in general.
10 The induced moduli morphism
In this section we discuss how the construction of Theorem A produces a morphism between moduli varieties. We retain the notation of the previous sections.
10.1 Moduli spaces on the curve and on the Jacobian
Fix integers $r\geq 2$ and $d$ with d>r(2g2)d>r(2g-2) and fix a line bundle $L_{0}\in Pic^{d}(C)$. Let
$M_{C}:=M_{C}^{s}(r,L_{0})$
denote the moduli variety of stable vector bundles on $C$ of rank $r$ and determinant $L_{0}$. It is a quasi-projective variety of dimension
$\dim M_{C}=(r^{2}-1)(g-1)$
(see [12, Ch. 4]).
Fix $m\geq g+1$ and set $H=m\Theta$. Let $\mathbf{v}_{U}$ denote the numerical invariants (rank and Chern character) of the Ulrich bundles $U$ constructed in Section 5 from bundles $V$ of rank $r$ and degree $d$. Let
$M_{A}:=M_{A}^{s}(\mathbf{v}_{U},H)$
be the moduli variety of $H$-stable vector bundles on $A=J(C)$ with invariants $\mathbf{v}_{U}$.
10.2 A universal family after passing to a cover
Since $M_{C}$ need not carry a universal bundle, we pass to a generically finite cover.
Lemma 10.1.
There exists a quasi-projective variety $T$ and a generically finite surjective morphism
$\pi:T\to M_{C}$
together with a vector bundle $\mathcal{V}$ on $C\times T$ such that for each $t\in T$,
$\mathcal{V}_{t}:=\mathcal{V}|_{C\times\{t\}}$
represents the stable bundle corresponding to $\pi(t)$.
Proof.
This follows from the GIT construction of $M_{C}$ as a quotient of an open subset of a Quot scheme; see [12, Ch. 4]. Restricting to the stable locus and passing to a suitable cover yields a universal bundle on $C\times T$. ∎
10.3 Relative Fourier–Mukai transform
Let $a:C\hookrightarrow A$ be the Abel–Jacobi embedding and $\mathcal{P}$ the normalized Poincaré bundle on $A\times\widehat{A}$, identified with $A$.
Define
$\mathcal{F}_{0}:=(a\times\mathrm{id}_{T})_{*}(\mathcal{V})\quad\text{on }A\times T,$
and set
$\mathcal{E}:=\Phi_{\mathcal{P}}(\mathcal{F}_{0})\quad\text{on }A\times T.$
By the slope hypothesis $\mu(\mathcal{V}_{t})>2g-2$ and the arguments of Section 3, WIT0 holds fiberwise. After shrinking $T$ to a dense open subset, we may assume:
  • (i)
    $\mathcal{E}$ is a vector bundle on $A\times T$,
  • (ii)
    $\mathcal{F}:=\mathcal{E}(\Theta)$ satisfies IT0 fiberwise (Section 4),
  • (iii)
    $rk(\mathcal{E}_{t})$ is constant and equals $h^{0}(C,\mathcal{V}_{t})$.
Openness of these properties follows from semicontinuity and cohomology and base change ([10, III.12], [7, §2]).
10.4 Relative evaluation and kernel construction
Fix the integer $N$ and general line bundles $\alpha_{1},\dots,\alpha_{N}\in Pic^{0}(A)$ as in Section 5.
Define on $A\times T$:
$\mathcal{H}_{k}:=p_{T*}(\mathcal{F}\otimes\alpha_{k}),\qquad\mathcal{G}:=\bigoplus_{k=1}^{N}p_{T}^{*}\mathcal{H}_{k}\otimes\alpha_{k}^{-1},$
where $p_{T}:A\times T\to T$.
Since $\mathcal{F}$ satisfies IT0, each $\mathcal{H}_{k}$ is locally free and base change holds.
The fiberwise evaluation maps
$G(V_{t})\twoheadrightarrow F(V_{t})$
constructed in Section 5 assemble into a morphism
$\mathrm{ev}:\ \mathcal{G}\longrightarrow\mathcal{F}$
which is surjective on a dense open subset of $T$ by continuous global generation ([7, Remark 2.2]).
After shrinking $T$, assume $\mathrm{ev}$ is surjective. Define the kernel
$0\longrightarrow\mathcal{M}\longrightarrow\mathcal{G}\longrightarrow\mathcal{F}\longrightarrow 0$
and set
$\mathcal{U}:=\mathcal{M}(m\Theta).$
Then $\mathcal{U}$ is flat over $T$ and its fiber over $t\in T$ is precisely the Ulrich bundle $U(V_{t})$ constructed in Section 5.
10.5 Stability in families
By Section 7, for general evaluation data the bundle $U(V_{t})$ is $\mu_{H}$-stable. Since stability is open in flat families ([12, Thm. 2.3.1]), there exists a dense open subset
$T^{\mathrm{st}}\subset T$
such that for all $t\in T^{\mathrm{st}}$, $\mathcal{U}_{t}$ is $H$-stable.
10.6 A morphism of moduli varieties
By the universal property of the coarse moduli space of stable bundles, the family $\mathcal{U}$ over $T^{\mathrm{st}}$ induces a morphism
$\Phi:\ T^{\mathrm{st}}\longrightarrow M_{A}^{s}(\mathbf{v}_{U},H),\qquad t\longmapsto[\mathcal{U}_{t}].$
This is a morphism of quasi-projective varieties.
10.7 Geometric consequences
Retaining all the notations in this section we state the geometric consequences as a theorem.
Theorem 10.2.
Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$. For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem A.
Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
$\pi:T\longrightarrow M_{C},$
and a dense open subset $T^{\mathrm{st}}\subset T$ such that the Ulrich construction is compatible with families and induces a morphism
$\Phi:T^{\mathrm{st}}\longrightarrow M_{A}.$
In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$.
Proof.
By Lemma 10.1, there exists a quasi-projective variety $T$ and a generically finite surjective morphism
$\pi:T\longrightarrow M_{C},$
together with a vector bundle $\mathcal{V}$ on $C\times T$ whose fiber $\mathcal{V}_{t}$ represents the point $\pi(t)\in M_{C}$.
Consider the family $(a\times id_{T})_{*}\mathcal{V}$ on $A\times T$ and apply the relative Fourier–Mukai transform with kernel the Poincaré bundle. The cohomology-and-base-change arguments of Sections 3 and 4, applied fiberwise using the hypothesis d>r(2g2)d>r(2g-2), show that after possibly shrinking $T$, the transform
$\mathcal{E}:=\Phi_{\mathcal{P}}\big((a\times id_{T})_{*}\mathcal{V}\big)$
is locally free on $A\times T$ and commutes with base change. Moreover, the twist $\mathcal{F}:=\mathcal{E}(\Theta)$ satisfies the $\mathrm{IT}_{0}$ property fiberwise (Section 4).
Since $\mathcal{F}$ is $\mathrm{IT}_{0}$ on each fiber, it is continuously globally generated in the sense of Pareschi–Popa. Fix line bundles $\alpha_{1},\dots,\alpha_{N}\in Pic^{0}(A)$ as in Section 5 and form the relative evaluation morphism
$\mathrm{ev}:\ \mathcal{G}:=\bigoplus_{k=1}^{N}p_{T}^{*}\!\big(p_{T*}(\mathcal{F}\otimes\alpha_{k})\big)\otimes\alpha_{k}^{-1}\longrightarrow\mathcal{F},$
where $p_{T}:A\times T\to T$ is the projection. By construction and upper semicontinuity of cohomology, after shrinking $T$ further if necessary, the morphism $\mathrm{ev}$ is surjective.
Let $\mathcal{M}:=\ker(\mathrm{ev})$ and define
$\mathcal{U}:=\mathcal{M}(m\Theta)\quad\text{on }A\times T.$
Then $\mathcal{U}$ is flat over $T$, and for every $t\in T$, the fiber $\mathcal{U}_{t}$ coincides with the Ulrich bundle on $(A,m\Theta)$ constructed from $\mathcal{V}_{t}$ in Theorem A (Section 5). In particular, each $\mathcal{U}_{t}$ is Ulrich with respect to $H=m\Theta$.
Let $T^{\mathrm{st}}\subset T$ be the locus where $\mathcal{U}_{t}$ is $H$-stable. Since slope-stability is an open property in flat families of coherent sheaves (see [12, Thm. 2.3.1]), the subset $T^{\mathrm{st}}$ is Zariski open. By the stability results of Section 8 (for general evaluation data and sufficiently positive $m$), this locus is nonempty; hence $T^{\mathrm{st}}$ is dense in $T$.
By the universal property of the coarse moduli space $M_{A}=M_{A}^{s}(v_{U},H)$, the family $\mathcal{U}|_{A\times T^{\mathrm{st}}}$ induces a morphism
$\Phi:T^{\mathrm{st}}\longrightarrow M_{A},\qquad t\longmapsto[\mathcal{U}_{t}],$
see [12, Cor. 4.6.6]. This proves the existence of $\pi$, $T^{\mathrm{st}}$, and $\Phi$ as stated.
Finally, since $\pi:T\to M_{C}$ is generically finite and dominant, one has
$\dim T=\dim M_{C}=(r^{2}-1)(g-1),$
and the same holds for the dense open subset $T^{\mathrm{st}}$. Therefore the image $\Phi(T^{\mathrm{st}})\subset M_{A}$ is swept out by a family of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$. In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles arising from stable bundles on $C$. ∎
Acknowledgements
I thank IISER Berhampur for its cohesive research environment and sustained support. I sincerely thank in advance to all readers of this arXiv version for kind comments/remarks/suggestions (if any).

### v2

@@
-arXiv:2603.02592v1[math.AG] 03 Mar 2026
+Ulrich Bundles on Jacobian Variety of a Curve
 Abstract
@@
 1 Introduction
-Ulrich bundles occupy a distinguished position in the geometry of projective embeddings.
-They may be viewed as extremal objects with respect to cohomological vanishing and linear resolutions.
-They play an important role in the study of syzygies, Chow forms, and derived categories [2, 6].
-A central problem is to determine which polarized varieties admit Ulrich bundles and to understand their geometric structure when they exist.
-The existence question has been resolved in several important cases, including hypersurfaces, complete intersections, and certain classes of surfaces.
-In particular, Beauville proved the existence of Ulrich bundles on abelian surfaces [1].In higher dimensions, however, systematic constructions remain comparatively sparse, even for classical varieties such as Jacobians.
-Beauville’s construction exploits special features of abelian surfaces and does not extend formally to higher-dimensional Jacobians.
-The present work provides a uniform construction in all dimensions, governed by slope positivity on the curve and Fourier–Mukai duality.
-In particular, the resulting Ulrich bundles arise functorially from stable bundles on $C$.
-Our construction produces positive-dimensional families in moduli space, linking the geometry of vector bundles on $C$ to Ulrich theory on $J(C)$.
-To the best of our knowledge, this provides the first general construction of Ulrich bundles on Jacobians in arbitrary dimension, together with a geometric description of the induced families in moduli.
-Let $C$ be a smooth complex projective curve of genus $g>1$, and let
-$A=J(C)$
-be its Jacobian endowed with its principal polarization $\Theta$.
-Although the geometry of $A$ is well understood, a uniform construction of Ulrich bundles on the polarized varieties $(A,m\Theta)$ has not previously been carried out in this generality.
-The purpose of this paper is to provide such a construction. Our starting point is a semistable vector bundle $V$ on $C$ satisfying the slope condition
-$\mu(V)>2g-2.$
-Let
-$a:C\hookrightarrow A$
-be the Abel–Jacobi embedding and let $\mathcal{P}$ denote the normalized Poincaré bundle on $A\times\widehat{A}$. We consider the Fourier–Mukai transform (see subsection 2.3)
-$E:=\Phi_{\mathcal{P}}(a_{*}V).$
-The key observation is that positivity of slope on the curve translates into strong cohomological control on the Jacobian.
-More precisely, the inequality $\mu(V)>2g-2$ implies uniform vanishing of $H^{1}(C,V\otimes M)$ for all $M\in Pic^{0}(C)$.
-Via derived base change, this yields the $\mathrm{WIT}_{0}$ property for $a_{*}V$ and shows that $E$ is locally free with explicitly computable rank.
-We then prove that the twist $E(\Theta)$ satisfies the stronger $\mathrm{IT}_{0}$ property by reducing cohomology on the Jacobian directly to cohomology on the curve.
-The $\mathrm{IT}_{0}$ property implies continuous global generation in the sense of Pareschi–Popa [7].
-A kernel construction then produces Ulrich bundles on $(A,m\Theta)$ for all sufficiently positive polarizations.
+Ulrich bundles occupy a central position in the interaction between algebraic geometry and commutative algebra.
+Originally introduced by Eisenbud and Schreyer in connection with linear determinantal representations and maximal Cohen–Macaulay modules [3].
+Ulrich bundles are vector bundles charaterized by strong cohomological vanishing properties with respect to given polarization.
+they play a central role in study of syzygies and linear resolutions of projective varieties and Chow forms of projective varieties.
+A fundamental conjecture of Eisenbud and Schreyer predicts that every polarized projective variety admits an Ulrich bundle.
+Despite substantial progress in special cases, the conjecture remains widely open.
+Ulrich bundles have been constructed for several classes of varieties, including many hypersurfaces and certain surfaces; see for example [1, 7].
+However, the existence problem remains largely unexplored for irregular varieties and higher-dimensional abelian varieties.
+In this paper we study the existence of Ulrich bundles on abelian varieties arising from Jacobians of smooth projective curves.
+Let $C$ be a smooth projective curve of genus $g\geq 2$, and let $A=J(C)$ denote its Jacobian equipped with the principal polarization $\Theta$.
+Our approach is based on Fourier–Mukai techniques and exploits the geometry of the Abel–Jacobi embedding $C\hookrightarrow J(C)$.
+Starting from a sufficiently positive line bundle $L$ on $C$, we consider the pushforward $a_{*}L$ along the Abel–Jacobi map and study its Fourier–Mukai transform.
+After a suitable twist by the principal polarization, we show that the resulting bundle on $J(C)$ satisfies strong regularity properties in the sense of Pareschi–Popa [8].
+Using the theory of continuously globally generated bundles on abelian varieties, we construct a kernel bundle whose twist produces an Ulrich bundle on $(J(C),m\Theta)$ for sufficiently large $m$.
+Our main result as follows.
 Theorem A (See Theorem 5.2).
@@
 Then for every integer $m\geq g+1$, the polarized Jacobian $(J(C),m\Theta)$ admits an Ulrich bundle constructed functorially from $V$.
-The rank of the resulting Ulrich bundle is determined by $h^{0}(C,V)$. Specializing to line bundles of degree >2g2>2g-2 yields an upper bound for the Ulrich complexity of $(J(C),m\Theta)$.
-Beyond existence, we analyze structural properties of the constructed bundles.
-Under natural stability assumptions on $V$, we prove that the resulting Ulrich bundles are slope-stable for large polarizations.
-Using Grothendieck–Riemann–Roch, we compute their Chern classes and show that all bundles arising in the construction have first Chern class proportional to the principal polarization $\Theta$.
-The construction is compatible with families and has geometric consequences at the level of moduli space.
-Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$.
-The above construction induces a natural morphism from a generically finite cover of $M_{C}$ to the moduli space of stable bundles on $A$ with the corresponding numerical invariants.
-This compatibility with families has strong geometric consequences. In particular, the Ulrich locus inside the moduli space of bundles on A is highly nontrivial:
+This construction verifies the Eisenbud–Schreyer conjecture for a large and geometrically natural class of abelian varieties, namely Jacobians of smooth projective curves endowed with sufficiently positive multiples of the principal polarization.
+To the best of our knowledge, this provides the first systematic construction of Ulrich bundles on Jacobians in arbitrary genus with respect to sufficiently positive polarization.
+It may be viewed as a higher-dimensional analogue of Beauville’s construction of Ulrich bundles on abelian surfaces [2].
+Beyond the existence statement, the construction has further geometric consequences.
+The Ulrich bundles obtained here arise naturally from vector bundles associated to the curve and vary in positive-dimensional families.
+More precisely, the construction induces families of stable bundles on the Jacobian arising from moduli spaces of vector bundles on the curve.
 Theorem B (See Theorem 10.2)
-Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$.
+Fix integers $r\geq 2,d>r(2g-2)$, and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$.
 For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem 5.2 Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
@@
 In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$.
-Thus the existence of Ulrich bundles on $(A,m\Theta)$ is governed by slope positivity on the curve and gives rise to a geometric link between vector bundles on $C$ and stable bundles on its Jacobian.
+These results illustrate a close relationship between the geometry of the curve and the moduli theory of vector bundles on its Jacobian.
+In particular, the construction connects Fourier–Mukai transforms, the geometry of Picard bundles, and the theory of Ulrich bundles.
+The techniques used here rely heavily on the modern theory of regularity on abelian varieties developed by Pareschi and Popa [8].
+Their results on $IT_{0}$ sheaves and continuously globally generated bundles provide the key tools for controlling the cohomological behaviour of the Fourier–Mukai transforms appearing in our construction.
+We also make use of Mukai’s theory of semihomogeneous vector bundles [16].
 Layout of the paper.
-Section 2 reviews background on slope stability and Fourier–Mukai transforms.
-Section 3 establishes the $\mathrm{WIT}_{0}$ property and computes the rank of the transform.
-Section 4 proves that $E(\Theta)$ satisfies $\mathrm{IT}_{0}$.
+In Section 2 we review some of the pre-requisites : Ulrich bundles, Abel-Jacobi map,Fourier–Mukai transforms, IT index properties, stable bundles and, semi-homogeneous bundles.
+We establish the $\mathrm{WIT}_{0}$ property and computes the rank of the transform in Section 3.
+In Section 4 we prove that $E(\Theta)$ satisfies $\mathrm{IT}_{0}$.
 Section 5 constructs Ulrich bundles and proves Theorem A.
-Section 6 discusses the numerical bound $m\geq g+1$.
+In Section 6 we discuss the numerical bound $m\geq g+1$.
 Section 7 studies slope stability.
-Section 8 computes Chern classes.
-Section 9 specializes to line bundles and derives an upper bound for Ulrich complexity.
-Section 10 constructs the moduli morphism and proves Theorem B.
+We compute Chern classes in Section 8.
+In Section 9 we specialize to line bundles and derive an upper bound for Ulrich complexity.
+We conclude Section 10 with the construction of the moduli morphism and proves the Theorem B.
+Throughout the paper we work over the field $\mathbb{C}$.
 2 Preliminaries
@@
 is a closed immersion. We identify $C$ with its image in $A$.
-A classical result (see [4, Ch. VII]) states that
+A classical result (see [5, Ch. VII]) states that
 $a^{*}\Theta\sim\text{a divisor of degree }g\text{ on }C.$
@@
     The sheaf $\mathcal{E}$ is called $\mu_{H}$-polystable if it is a direct sum of $\mu_{H}$-stable sheaves all having the same slope.
-(See [12, Ch. 1] )
+(See [13, Ch. 1] )
 In particular, a direct sum of $\mu_{H}$-stable sheaves of identical slope is $\mu_{H}$-polystable.
@@
 Semihomogeneous bundles were introduced and studied by Mukai [16]. They enjoy the following properties:
-  • A semihomogeneous bundle is simple if and only if it is stable with respect to any polarization [15, Prop. 6.16].
-  • If $\mathcal{E}$ is semihomogeneous, then its Chern classes are proportional to powers of $c_{1}(\mathcal{E})$ in the Néron–Severi group [15, Thm. 5.8].
+  • A semihomogeneous bundle is simple if and only if it is stable with respect to any polarization [16, Prop. 6.16].
+  • If $\mathcal{E}$ is semihomogeneous, then its Chern classes are proportional to powers of $c_{1}(\mathcal{E})$ in the Néron–Severi group [16, Thm. 5.8].
 3 Cohomological Vanishing and the WIT0 Property
@@
 For $k\geq 1$, the line bundle $\mathcal{O}_{A}(k\Theta)$ is ample.
-On an abelian variety, twisting an IT0 sheaf by an ample line bundle preserves the vanishing of higher cohomology (see [7, Thm.
+On an abelian variety, twisting an IT0 sheaf by an ample line bundle preserves the vanishing of higher cohomology (see [8, Thm.
 2.3]).
@@
 $H^{i}(A,G(H-pH))=0\quad\forall i>0.$
+Infact, taking each $\alpha_{1},\alpha_{2},.....,\alpha_{N}$ non-trivial line bundles, we can conclude that
+$H^{0}A,G(H-pH))=0\quad\forall p\geq 1$
+Thus we conclude
+$H^{i}(A,G(H-pH))=0\quad\forall i\geq 0,\forall p\geq 1$
 Vanishing for F(HpH)F(H-pH).
@@
 $Hom_{A}(a_{*}V,a_{*}V)\;\cong\;Hom_{C}(V,V).$
-Because $V$ is stable on a curve, $End_{C}(V)=\mathbb{C}$ (see [12, Ch. 1]), hence $Hom_{A}(a_{*}V,a_{*}V)=\mathbb{C}$.
+Because $V$ is stable on a curve, $End_{C}(V)=\mathbb{C}$ (see [13, Ch. 1]), hence $Hom_{A}(a_{*}V,a_{*}V)=\mathbb{C}$.
 Now we apply the Fourier–Mukai equivalence $\Phi_{\mathcal{P}}:D^{b}(A)\to D^{b}(A)$. Since $\Phi_{\mathcal{P}}$ is an equivalence, it preserves $Hom$ in the derived category:
@@
 The sheaf $E=\Phi_{\mathcal{P}}(a_{*}V)$ is a Picard sheaf in the sense of Mukai (Fourier–Mukai transform of a sheaf pushed forward from a curve inside its Jacobian).
-Mukai proves that Picard sheaves on abelian varieties are semihomogeneous; see [3, §3–§4] (and also the exposition [14, Ch.
+Mukai proves that Picard sheaves on abelian varieties are semihomogeneous; see [4, §3–§4] (and also the exposition [15, Ch.
 6]).
 Thus $E$ is semihomogeneous, and twisting by a line bundle preserves semihomogeneity, so $F=E(\Theta)$ is semihomogeneous as well.
-Finally, by Mukai’s stability criterion, a simple semihomogeneous vector bundle on an abelian variety is $\mu_{H}$-stable with respect to every polarization $H$; see [5, Prop.
-6.16] (or [14, Prop.
+Finally, by Mukai’s stability criterion, a simple semihomogeneous vector bundle on an abelian variety is $\mu_{H}$-stable with respect to every polarization $H$; see [6, Prop.
+6.16] (or [15, Prop.
 6.16]).
@@
 $c_{1}(\mathcal{P})=\sum_{i}\alpha_{i}\otimes\beta_{i},$
-with $\{\alpha_{i}\}$ a basis of $H^{1}(C)$ and $\{\beta_{i}\}$ the dual basis of $H^{1}(A)$ (see [14, Ch. 6]).
+with $\{\alpha_{i}\}$ a basis of $H^{1}(C)$ and $\{\beta_{i}\}$ the dual basis of $H^{1}(A)$ (see [15, Ch. 6]).
 Only terms of total codimension two on $C\times A$ contribute to the first Chern class of $E$ after pushing forward.
@@
 $r\cdot c_{1}(\mathcal{P})^{2}.$
-A classical computation (see [14, Ch. 6] and [5, Prop. 6.16]) shows that
+A classical computation (see [15, Ch. 6] and [6, Prop. 6.16]) shows that
 $p_{2*}\!\left(c_{1}(\mathcal{P})^{2}\right)=2\,\Theta.$
@@
 $E=\Phi_{\mathcal{P}}(a_{*}L),\qquad F=E(\Theta).$
-By Theorem 4.1, $F$ satisfies IT0. In particular, $F$ is continuously globally generated (CGG) in the sense of Pareschi–Popa (see [7, Remark 2.2(b)]).
+By Theorem 4.1, $F$ satisfies IT0. In particular, $F$ is continuously globally generated (CGG) in the sense of Pareschi–Popa (see [8, Remark 2.2(b)]).
 Hence there exist finitely many line bundles
@@
 $\dim M_{C}=(r^{2}-1)(g-1)$
-(see [12, Ch. 4]).
+(see [13, Ch. 4]).
 Fix $m\geq g+1$ and set $H=m\Theta$.
@@
 Proof.
-This follows from the GIT construction of $M_{C}$ as a quotient of an open subset of a Quot scheme; see [12, Ch.
+This follows from the GIT construction of $M_{C}$ as a quotient of an open subset of a Quot scheme; see [13, Ch.
 4].
@@
     $rk(\mathcal{E}_{t})$ is constant and equals $h^{0}(C,\mathcal{V}_{t})$.
-Openness of these properties follows from semicontinuity and cohomology and base change ([10, III.12], [7, §2]).
+Openness of these properties follows from semicontinuity and cohomology and base change ([11, III.12], [8, §2]).
 10.4 Relative evaluation and kernel construction
@@
 $\mathrm{ev}:\ \mathcal{G}\longrightarrow\mathcal{F}$
-which is surjective on a dense open subset of $T$ by continuous global generation ([7, Remark 2.2]).
+which is surjective on a dense open subset of $T$ by continuous global generation ([8, Remark 2.2]).
 After shrinking $T$, assume $\mathrm{ev}$ is surjective. Define the kernel
@@
 10.5 Stability in families
-By Section 7, for general evaluation data the bundle $U(V_{t})$ is $\mu_{H}$-stable. Since stability is open in flat families ([12, Thm. 2.3.1]), there exists a dense open subset
+By Section 7, for general evaluation data the bundle $U(V_{t})$ is $\mu_{H}$-stable. Since stability is open in flat families ([13, Thm. 2.3.1]), there exists a dense open subset
 $T^{\mathrm{st}}\subset T$
@@
 Let $T^{\mathrm{st}}\subset T$ be the locus where $\mathcal{U}_{t}$ is $H$-stable.
-Since slope-stability is an open property in flat families of coherent sheaves (see [12, Thm.
+Since slope-stability is an open property in flat families of coherent sheaves (see [13, Thm.
 2.3.1]), the subset $T^{\mathrm{st}}$ is Zariski open.
@@
 $\Phi:T^{\mathrm{st}}\longrightarrow M_{A},\qquad t\longmapsto[\mathcal{U}_{t}],$
-see [12, Cor. 4.6.6]. This proves the existence of $\pi$, $T^{\mathrm{st}}$, and $\Phi$ as stated.
+see [13, Cor. 4.6.6]. This proves the existence of $\pi$, $T^{\mathrm{st}}$, and $\Phi$ as stated.
 Finally, since $\pi:T\to M_{C}$ is generically finite and dominant, one has

### v3

@@
-Ulrich Bundles on Jacobian Variety of a Curve
+Pabitra Barik
 Abstract
@@
 We prove that $E(\Theta)$ satisfies the $\mathrm{IT}_{0}$ property.
-As a consequence, for every $m\geq g+1$, the polarized Jacobian $(A,m\Theta)$ admits Ulrich bundles constructed functorially from $V$.
-Further we analyze stability and Chern classes of the resulting bundles.
-We show that the construction induces a natural morphism from the generically finite cover of moduli space of stable bundles on $C$ to the moduli space of stable bundles on $A$, producing positive-dimensional families of stable Ulrich bundles.
 1 Introduction
-Ulrich bundles occupy a central position in the interaction between algebraic geometry and commutative algebra.
-Originally introduced by Eisenbud and Schreyer in connection with linear determinantal representations and maximal Cohen–Macaulay modules [3].
-Ulrich bundles are vector bundles charaterized by strong cohomological vanishing properties with respect to given polarization.
-they play a central role in study of syzygies and linear resolutions of projective varieties and Chow forms of projective varieties.
-A fundamental conjecture of Eisenbud and Schreyer predicts that every polarized projective variety admits an Ulrich bundle.
-Despite substantial progress in special cases, the conjecture remains widely open.
-Ulrich bundles have been constructed for several classes of varieties, including many hypersurfaces and certain surfaces; see for example [1, 7].
-However, the existence problem remains largely unexplored for irregular varieties and higher-dimensional abelian varieties.
-In this paper we study the existence of Ulrich bundles on abelian varieties arising from Jacobians of smooth projective curves.
 Let $C$ be a smooth projective curve of genus $g\geq 2$, and let $A=J(C)$ denote its Jacobian equipped with the principal polarization $\Theta$.
 Our approach is based on Fourier–Mukai techniques and exploits the geometry of the Abel–Jacobi embedding $C\hookrightarrow J(C)$.
-Starting from a sufficiently positive line bundle $L$ on $C$, we consider the pushforward $a_{*}L$ along the Abel–Jacobi map and study its Fourier–Mukai transform.
+Starting from a sufficiently positive semistable vector bundle bundle $V$ on $C$, we consider the pushforward $a_{*}V$ along the Abel–Jacobi map and study its Fourier–Mukai transform.
 After a suitable twist by the principal polarization, we show that the resulting bundle on $J(C)$ satisfies strong regularity properties in the sense of Pareschi–Popa [4].
-Using the theory of continuously globally generated bundles on abelian varieties, we construct a kernel bundle whose twist produces an Ulrich bundle on $(J(C),m\Theta)$ for sufficiently large $m$.
-Our main result as follows.
-Theorem A (See Theorem 5.2).
-Let $C$ be a smooth projective curve of genus $g>1$ and let $V$ be a semistable vector bundle on $C$ with $\mu(V)>2g-2$.
-Then for every integer $m\geq g+1$, the polarized Jacobian $(J(C),m\Theta)$ admits an Ulrich bundle constructed functorially from $V$.
-This construction verifies the Eisenbud–Schreyer conjecture for a large and geometrically natural class of abelian varieties, namely Jacobians of smooth projective curves endowed with sufficiently positive multiples of the principal polarization.
-To the best of our knowledge, this provides the first systematic construction of Ulrich bundles on Jacobians in arbitrary genus with respect to sufficiently positive polarization.
-It may be viewed as a higher-dimensional analogue of Beauville’s construction of Ulrich bundles on abelian surfaces [2].
-Beyond the existence statement, the construction has further geometric consequences.
-The Ulrich bundles obtained here arise naturally from vector bundles associated to the curve and vary in positive-dimensional families.
-More precisely, the construction induces families of stable bundles on the Jacobian arising from moduli spaces of vector bundles on the curve.
-Theorem B (See Theorem 10.2)
-Fix integers $r\geq 2,d>r(2g-2)$, and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$.
-For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem 5.2 Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
-$\pi:T\longrightarrow M_{C},$
-and a dense open subset $T^{\mathrm{st}}\subset T$ such that the Ulrich construction is compatible with families and induces a morphism
-$\Phi:T^{\mathrm{st}}\longrightarrow M_{A}.$
-In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$.
-These results illustrate a close relationship between the geometry of the curve and the moduli theory of vector bundles on its Jacobian.
-In particular, the construction connects Fourier–Mukai transforms, the geometry of Picard bundles, and the theory of Ulrich bundles.
-The techniques used here rely heavily on the modern theory of regularity on abelian varieties developed by Pareschi and Popa [8].
-Their results on $IT_{0}$ sheaves and continuously globally generated bundles provide the key tools for controlling the cohomological behaviour of the Fourier–Mukai transforms appearing in our construction.
-We also make use of Mukai’s theory of semihomogeneous vector bundles [16].
-Layout of the paper.
-In Section 2 we review some of the pre-requisites : Ulrich bundles, Abel-Jacobi map,Fourier–Mukai transforms, IT index properties, stable bundles and, semi-homogeneous bundles.
-We establish the $\mathrm{WIT}_{0}$ property and computes the rank of the transform in Section 3.
-In Section 4 we prove that $E(\Theta)$ satisfies $\mathrm{IT}_{0}$.
-Section 5 constructs Ulrich bundles and proves Theorem A.
-In Section 6 we discuss the numerical bound $m\geq g+1$.
-Section 7 studies slope stability.
-We compute Chern classes in Section 8.
-In Section 9 we specialize to line bundles and derive an upper bound for Ulrich complexity.
-We conclude Section 10 with the construction of the moduli morphism and proves the Theorem B.
-Throughout the paper we work over the field $\mathbb{C}$.
-2 Preliminaries
-2.1 Ulrich bundles and Ulrich complexity
-Let $(X,H)$ be a polarized smooth projective variety of dimension $n$, where $H$ is an ample divisor on $X$. A vector bundle $E$ on $X$ is called an Ulrich bundle with respect to $H$ if
-$H^{i}(X,E(-jH))=0\qquad\text{for all }i\text{ and }1\leq j\leq n.$
-In this paper we will consider Ulrich bundles on polarized Jacobians
-$(A,m\Theta),$
-where $A=J(C)$ is the Jacobian of a smooth complex projective curve $C$ of genus $g>1$ and $\Theta$ denotes the principal polarization.
-Ulrich complexity. For a polarized variety $(X,H)$ admitting Ulrich bundles, we define the Ulrich complexity to be
-$\mathrm{uc}(X,H):=\min\{rk(E)\;:\;E\text{ is an Ulrich bundle on }(X,H)\}.$
-In general, determining $\mathrm{uc}(X,H)$ is a subtle problem.
-In the present work, we construct explicit Ulrich bundles on $(J(C),m\Theta)$ and derive upper bounds for $\mathrm{uc}(J(C),m\Theta)$ in terms of the genus $g$ and the degree of the initial bundle on $C$.
+We recall some preliminaries.
 1.1 The Abel–Jacobi embedding
@@
 $a:C\longrightarrow A,\qquad x\longmapsto\mathcal{O}_{C}(x-x_{0}),$
-is a closed immersion. We identify $C$ with its image in $A$.
-A classical result (see [5, Ch. VII]) states that
+is a closed immersion. We identify $C$ with its image in AA.
+A classical result (see [2, Ch. VII]) states that
 $a^{*}\Theta\sim\text{a divisor of degree }g\text{ on }C.$
@@
 1.2 Fourier–Mukai transform on the Jacobian
-Let $\widehat{A}:=Pic^{0}(A)$ be the dual abelian variety. Since $A$ is principally polarized, we identify
+Let $\widehat{A}:=Pic^{0}(A)$ be the dual abelian variety. Since AA is principally polarized, we identify
 $A\cong\widehat{A}$
@@
 1.3 WIT and IT conditions
-Let $\mathcal{F}$ be a coherent sheaf on $A$.
-Definition 2.1.
-We say that $\mathcal{F}$ satisfies WITi (Weak Index Theorem of index $i$) if
+Let $\mathcal{F}$ be a coherent sheaf on AA.
+Definition $1.$
+We say that $\mathcal{F}$ satisfies WITi (Weak Index Theorem of index ii) if
 $R^{j}\Phi_{\mathcal{P}}(\mathcal{F})=0\quad\text{for all }j\neq i.$
@@
 Definition 1.2.
-A coherent sheaf $\mathcal{E}$ on $A$ satisfies IT0 (Index Theorem with index 0) if
+A coherent sheaf $\mathcal{E}$ on AA satisfies IT0 (Index Theorem with index 0) if
 $H^{i}(A,\mathcal{E}\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in Pic^{0}(A).$
@@
 1.4 Base change for Fourier–Mukai
-Let $\mathcal{F}$ be a coherent sheaf on $A$. For $\alpha\in Pic^{0}(A)$, derived base change gives
+Let $\mathcal{F}$ be a coherent sheaf on AA. For $\alpha\in Pic^{0}(A)$, derived base change gives
 $\Phi_{\mathcal{P}}(\mathcal{F})\otimes^{\mathbf{L}}k(\alpha)\simeq R\Gamma(A,\mathcal{F}\otimes\mathcal{P}_{\alpha}).$
@@
     The sheaf $\mathcal{E}$ is called $\mu_{H}$-polystable if it is a direct sum of $\mu_{H}$-stable sheaves all having the same slope.
-(See [13, Ch. 1] )
+(See [7, Ch. 1] )
 In particular, a direct sum of $\mu_{H}$-stable sheaves of identical slope is $\mu_{H}$-polystable.
@@
 for any vector bundle $V$ on $C$, and semistability is independent of a choice of polarization. Semistability is preserved under tensoring with line bundles of degree zero.
-2.7 Semihomogeneous bundles on abelian varieties
-Let $A$ be an abelian variety and denote by $t_{x}:A\to A$ the translation by a point $x\in A$.
-Definition 2.4.
-A vector bundle $\mathcal{E}$ on $A$ is called semihomogeneous if for every point $x\in A$ there exists a line bundle $\alpha_{x}\in Pic^{0}(A)$ such that
-$t_{x}^{*}\mathcal{E}\cong\mathcal{E}\otimes\alpha_{x}.$
-Semihomogeneous bundles were introduced and studied by Mukai [16]. They enjoy the following properties:
-  • A semihomogeneous bundle is simple if and only if it is stable with respect to any polarization [16, Prop. 6.16].
-  • If $\mathcal{E}$ is semihomogeneous, then its Chern classes are proportional to powers of $c_{1}(\mathcal{E})$ in the Néron–Severi group [16, Thm. 5.8].
 2 Cohomological Vanishing and the WIT0 Property
@@
 The claim follows. ∎
-Remark 3.2.
+Remark $2.$
 The slope inequality $\mu(V)>2g-2$ is sharp for this argument. It ensures that $V$ admits no nonzero morphism to $\omega_{C}$, which is precisely the obstruction to the vanishing of $H^{1}$.
@@
   2. 2.
-    $E$ is a locally free sheaf on $A$;
+    $E$ is a locally free sheaf on AA;
   3. 3.
@@
 $h^{0}(C,V)=\deg(V)+\operatorname{rk}(V)(1-g).$
-The result follows from Proposition $3.$ ∎
+The result follows from Proposition 2.3. ∎
 The results of this section show that positivity of slope on the curve translates into strong structural properties of the Fourier–Mukai transform on the Jacobian.
@@
 4.4 Vanishing of higher cohomology
-Since $C$ has dimension $1$, we have
+Since $C$ has dimension 11, we have
 $H^{i}(C,-)=0\quad\text{for }i\geq 2.$
@@
 The vanishing for $i=1$ follows from the slope computation and Lemma 3.1. The vanishing for $i\geq 2$ follows from the fact that cohomology on the curve vanishes in degrees $\geq 2$. ∎
-5 Construction of Ulrich Bundles
-In this section we complete the proof of Theorem A by constructing Ulrich bundles on $(A,m\Theta)$ from the IT0 property established in Section 4.
-Throughout this section we assume that $V$ is semistable with $\mu(V)>2g-2$, and we set
-$E:=\Phi_{\mathcal{P}}(a_{*}V).$
-By Theorem 4.1,
-$F:=E(\Theta)$
-satisfies IT0.
-5.1 Positivity and higher cohomology vanishing
-We first record a simple consequence of the IT0 property.
-Lemma 5.1.
-Let $F=E(\Theta)$ be as above. If $F$ satisfies IT0, then for every integer $k\geq 1$,
-$H^{i}(A,F(k\Theta)\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in\mathrm{Pic}^{0}(A).$
-In particular,
-$H^{i}(A,E((k+1)\Theta))=0\quad\forall i>0,\ k\geq 1.$
-Proof.
-Since $F$ satisfies IT0, we have
-$H^{i}(A,F\otimes\beta)=0\quad\forall i>0,\ \forall\beta\in\mathrm{Pic}^{0}(A).$
-For $k\geq 1$, the line bundle $\mathcal{O}_{A}(k\Theta)$ is ample.
-On an abelian variety, twisting an IT0 sheaf by an ample line bundle preserves the vanishing of higher cohomology (see [8, Thm.
-2.3]).
-Hence
-$H^{i}(A,F(k\Theta)\otimes\alpha)=0\quad\forall i>0,\ \forall\alpha\in\mathrm{Pic}^{0}(A).$
-Taking $\alpha=\mathcal{O}_{A}$ gives the second statement. ∎
-5.2 Continuous global generation
-A fundamental consequence of the IT0 property is continuous global generation.
-By [8, 9], a locally free sheaf satisfying IT0 is continuously globally generated. Hence there exist finitely many
-$\alpha_{1},\dots,\alpha_{N}\in\mathrm{Pic}^{0}(A)$
-such that the evaluation morphism
-$\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}\longrightarrow F$
-is surjective.
-Let
-$0\longrightarrow M\longrightarrow G\longrightarrow F\longrightarrow 0$
-be the resulting exact sequence, where
-$G:=\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}.$
-5.3 Twisting and Ulrich vanishing
-Let $H=m\Theta$ with $m\geq g+1$, and define
-$U:=M(H).$
-We verify that $U$ is Ulrich with respect to $H$.
-The vector bundle $U$ on $(A,H)$ is Ulrich if
-$H^{i}(A,U(-pH))=0\quad\forall i,\ 1\leq p\leq g.$
-For $1\leq p\leq g$, twist the above evaluation sequence by (HpH)(H-pH):
-$0\to M(H-pH)\to G(H-pH)\to F(H-pH)\to 0.$
-Since $U=M(H)$, we have
-U(pH)=M(HpH).U(-pH)=M(H-pH).
-We analyze the two terms.
-Vanishing for G(HpH)G(H-pH).
-Each summand of $G$ is of the form $\alpha_{k}^{-1}$. Hence
-$G(H-pH)=\bigoplus_{k}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}((m-p)\Theta).$
-If $1\leq p\leq g$ and $m\geq g+1$, then $(m-p)\Theta$ is ample. On an abelian variety, an ample line bundle satisfies
-$H^{i}(A,\mathcal{O}_{A}((m-p)\Theta)\otimes\alpha)=0\quad\forall i>0.$
-Thus
-$H^{i}(A,G(H-pH))=0\quad\forall i>0.$
-Infact, taking each $\alpha_{1},\alpha_{2},.....,\alpha_{N}$ non-trivial line bundles, we can conclude that
-$H^{0}A,G(H-pH))=0\quad\forall p\geq 1$
-Thus we conclude
-$H^{i}(A,G(H-pH))=0\quad\forall i\geq 0,\forall p\geq 1$
-Vanishing for F(HpH)F(H-pH).
-We compute
-$F(H-pH)=E(\Theta+(m-p)\Theta)=E((m-p+1)\Theta).$
-Since $m\geq g+1$ and $1\leq p\leq g$, we have
-$m-p+1\geq 2.$
-Thus F(HpH)F(H-pH) is a positive twist of $F$. By Lemma 5.1,
-$H^{i}(A,F(H-pH))=0\quad\forall i>0.$
-5.4 Conclusion
-The long exact sequence in cohomology associated to the twisted evaluation sequence yields
-$H^{i}(A,U(-pH))=0\quad\forall i,\ 1\leq p\leq g.$
-Hence $U$ is an Ulrich bundle on $(A,m\Theta)$.
-Hence we conclude the main theorem
-Theorem 5.2.
-Let $C$ be a smooth projective curve of genus $g>1$ and let $V$ be a semistable vector bundle on $C$ with $\mu(V)>2g-2$.
-Then for every integer $m\geq g+1$, the polarized Jacobian $(J(C),m\Theta)$ admits an Ulrich bundle constructed functorially from $V$.
-Remark 5.3.
-We do not claim optimality of the bound $m\geq g+1$. The construction is functorial in $V$, and the rank of the resulting Ulrich bundle depends explicitly on $h^{0}(C,V)$.
-6 On the bound $m\geq g+1$
-In Theorem 5.2 we obtained Ulrich bundles on $(A,m\Theta)$ for all integers $m\geq g+1$. We briefly explain why this numerical condition naturally arises from the construction.
-Recall that $U=M(m\Theta)$, where
-$0\to M\to G\to F\to 0,\qquad F=E(\Theta),$
-and $F$ satisfies $\mathrm{IT}_{0}$. To verify the Ulrich condition, one must prove that
-$H^{i}(A,U(-pH))=0\quad\text{for all }i\text{ and }1\leq p\leq g,$
-where $H=m\Theta$.
-Since
-$U(-pH)=M((m-p)\Theta),$
-we twist the evaluation sequence by $(m-p)\Theta$. The argument in Section 5 uses two facts:
-  1. (i)
-    $(m-p)\Theta$ must be ample to ensure vanishing for the bundle $G$;
-  2. (ii)
-    the twist $E((m-p+1)\Theta)$ must remain a positive twist of $E(\Theta)$ so that Lemma 5.1 applies.
-Because $\dim A=g$, the most restrictive case is $p=g$. Thus we must have
-$m-g+1\geq 2,$
-which is equivalent to
-$m\geq g+1.$
-If $m=g$, then the extremal term involves $E(\Theta)$ itself, and the argument does not guarantee the required vanishing.
-If $m<g$, some twists $(m-p)\Theta$ cease to be ample, and the cohomological vanishing used in Section 5 breaks down.
-Therefore the condition $m\geq g+1$ is intrinsic to the present method. We do not claim optimality of this bound; the existence of Ulrich bundles for smaller values of $m$ remains open.
-7 Stability of the Ulrich Bundle
-In this section we study slope stability of the Ulrich bundles constructed in Section 5.
-Throughout, let $C$ be a smooth projective curve of genus $g>1$, let $A=J(C)$ be its Jacobian with principal polarization $\Theta$, and let $V$ be a stable vector bundle on $C$ with $\mu(V)>2g-2$. Set
-$E:=\Phi_{\mathcal{P}}(a_{*}V),\qquad F:=E(\Theta).$
-By Theorem 4.1, $F$ satisfies IT0.
-Stability of $F$
-We first see that Fourier–Mukai transforms of stable bundles supported on a curve often yield simple semihomogeneous bundles on the Jacobian.
-Proposition 7.1.
-Assume that $V$ is stable on $C$. Then $F=E(\Theta)$ is a simple semihomogeneous vector bundle on $A$. In particular, $F$ is $\mu_{\Theta}$-stable.
-Proof.
-We first prove simplicity. Since $a:C\hookrightarrow A$ is a closed immersion, the pushforward functor $a_{*}$ is exact and fully faithful on coherent sheaves supported on $C$; in particular,
-$Hom_{A}(a_{*}V,a_{*}V)\;\cong\;Hom_{C}(V,V).$
-Because $V$ is stable on a curve, $End_{C}(V)=\mathbb{C}$ (see [13, Ch. 1]), hence $Hom_{A}(a_{*}V,a_{*}V)=\mathbb{C}$.
-Now we apply the Fourier–Mukai equivalence $\Phi_{\mathcal{P}}:D^{b}(A)\to D^{b}(A)$. Since $\Phi_{\mathcal{P}}$ is an equivalence, it preserves $Hom$ in the derived category:
-$Hom_{D^{b}(A)}(a_{*}V,a_{*}V)\;\cong\;Hom_{D^{b}(A)}(\Phi_{\mathcal{P}}(a_{*}V),\Phi_{\mathcal{P}}(a_{*}V)).$
-By Proposition 3.3 in this paper, $E=\Phi_{\mathcal{P}}(a_{*}V)$ is a vector bundle, hence $Hom_{D^{b}(A)}(E,E)=Hom_{A}(E,E)$.
-Combining with the previous paragraph gives $End_{A}(E)=\mathbb{C}$, i.e.
-$E$ is simple; therefore $F=E(\Theta)$ is also simple.
-Next we explain semihomogeneity.
-The sheaf $E=\Phi_{\mathcal{P}}(a_{*}V)$ is a Picard sheaf in the sense of Mukai (Fourier–Mukai transform of a sheaf pushed forward from a curve inside its Jacobian).
-Mukai proves that Picard sheaves on abelian varieties are semihomogeneous; see [4, §3–§4] (and also the exposition [15, Ch.
-6]).
-Thus $E$ is semihomogeneous, and twisting by a line bundle preserves semihomogeneity, so $F=E(\Theta)$ is semihomogeneous as well.
-Finally, by Mukai’s stability criterion, a simple semihomogeneous vector bundle on an abelian variety is $\mu_{H}$-stable with respect to every polarization $H$; see [6, Prop.
-6.16] (or [15, Prop.
-6.16]).
-Taking $H=\Theta$ yields that $F$ is $\mu_{\Theta}$-stable.
-∎
-Slope of the evaluation bundle
-We recall from Section 5 that, since $F$ satisfies IT0, it is continuously globally generated. Thus there exist finitely many $\alpha_{1},\dots,\alpha_{N}\in\mathrm{Pic}^{0}(A)$ and a surjection
-$\varphi:\;G:=\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}\longrightarrow F.$
-Let
-$0\longrightarrow M\longrightarrow G\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}F\longrightarrow 0$
-be the corresponding exact sequence.
-Lemma 7.2.
-$G$ is $\mu_{\Theta}$-polystable of slope 0.
-Proof.
-Each summand $\alpha_{k}^{-1}$ lies in $\mathrm{Pic}^{0}(A)$ and thus has trivial first Chern class.
-Hence $c_{1}(\alpha_{k}^{-1})=0$, so $\mu_{\Theta}(\alpha_{k}^{-1})=0$.
-Being a direct sum of degree-zero line bundles, $G$ is polystable of slope 0.
-∎
-Since $F$ is $\mu_{\Theta}$-stable and $c_{1}(F)$ is proportional to $\Theta$, we have
-$\mu_{\Theta}(F)>0.$
-Stability of the kernel
-We now prove stability of the kernel for a general choice of evaluation map.
-Proposition 7.3.
-Assume $F$ is $\mu_{\Theta}$-stable. For a general surjection $\varphi:G\twoheadrightarrow F$ as above, the kernel $M=\ker(\varphi)$ is $\mu_{\Theta}$-stable.
-Proof.
-Suppose that $M$ is not stable. Then there exists a proper saturated subsheaf $K\subset M$ with
-$\mu_{\Theta}(K)\geq\mu_{\Theta}(M).$
-Composing $K\hookrightarrow M\hookrightarrow G$, we obtain a nonzero subsheaf of $G$. Since $G$ is polystable of slope 0, any subsheaf has slope $\leq 0$.
-On the other hand, from the exact sequence we compute
-$\mu_{\Theta}(M)=\frac{rk(G)\mu_{\Theta}(G)-rk(F)\mu_{\Theta}(F)}{rk(M)}=-\frac{rk(F)}{rk(M)}\,\mu_{\Theta}(F)<0.$
-Thus any destabilizing $K$ would satisfy
-$\mu_{\Theta}(K)\geq\mu_{\Theta}(M)>-\mu_{\Theta}(F).$
-A standard parameter-count argument in $\operatorname{Hom}(G,F)$ shows that the existence of such a $K$ forces $\varphi$ to lie in a proper closed subset of $\operatorname{Hom}(G,F)$.
-Hence for a general choice of $\varphi$, no such destabilizing $K$ exists.
-Therefore $M$ is $\mu_{\Theta}$-stable.
-∎
-Stability of the Ulrich bundle
-Finally, let $H=m\Theta$ with $m\geq g+1$ and define
-$U:=M(H).$
-Theorem 7.4.
-Under the above assumptions, for a general choice of evaluation map and for all $m\gg 0$, the Ulrich bundle $U$ on $(A,m\Theta)$ is $\mu_{m\Theta}$-stable.
-Proof.
-Twisting by a line bundle preserves slope stability. Hence $U$ is $\mu_{m\Theta}$-stable if and only if $M$ is $\mu_{m\Theta}$-stable.
-Stability is an open condition with respect to the polarization ray $\mathbb{R}_{>0}\Theta$. Since $M$ is $\mu_{\Theta}$-stable for general $\varphi$, it remains stable for $m\gg 0$. ∎
-8 Chern Class
-In this section we compute the Chern character of the Fourier–Mukai bundle
-$E=\Phi_{\mathcal{P}}(a_{*}V)$
-and determine the first Chern classes of the bundles appearing in our Ulrich construction.
-These computations clarify the numerical structure of the construction and will be useful for slope considerations.
-Geometric setup
-Let $C$ be a smooth projective curve of genus $g>1$, let $A=J(C)$ be its Jacobian with principal polarization $\Theta$, and let $a:C\hookrightarrow A$ denote the Abel–Jacobi embedding.
-Let $V$ be a vector bundle on $C$ of rank $r$ and degree $d$. We assume throughout that $\mu(V)>2g-2$, so that the Fourier–Mukai transform
-$E=\Phi_{\mathcal{P}}(a_{*}V)$
-is locally free (Section 3).
-Recall that
-$E=Rp_{2*}\!\left(p_{1}^{*}a_{*}V\otimes\mathcal{P}\right),$
-where $\mathcal{P}$ is the normalized Poincaré bundle on $A\times A$ and $p_{1},p_{2}$ are the projections.
-Since $a_{*}V$ is supported on the embedded curve $C\subset A$, this expression reduces to a pushforward along
-$p_{2}:C\times A\longrightarrow A.$
-Applying Grothendieck–Riemann–Roch
-We apply Grothendieck–Riemann–Roch to the morphism $p_{2}$. It gives
-$\mathrm{ch}(E)=p_{2*}\left(\mathrm{ch}(p_{1}^{*}V)\cdot\mathrm{ch}(\mathcal{P}|_{C\times A})\cdot\mathrm{td}(C)\right).$
-Since $C$ has dimension one, the Chern character and Todd class simplify considerably:
-$\mathrm{ch}(V)=r+d\,[\mathrm{pt}],\qquad\mathrm{td}(C)=1+(1-g)\,[\mathrm{pt}].$
-Thus
-$\mathrm{ch}(V)\cdot\mathrm{td}(C)=r+(d+r(1-g))[\mathrm{pt}].$
-The Poincaré bundle satisfies
-$\mathrm{ch}(\mathcal{P})=1+c_{1}(\mathcal{P})+\tfrac{1}{2}c_{1}(\mathcal{P})^{2}+\cdots,$
-where
-$c_{1}(\mathcal{P})=\sum_{i}\alpha_{i}\otimes\beta_{i},$
-with $\{\alpha_{i}\}$ a basis of $H^{1}(C)$ and $\{\beta_{i}\}$ the dual basis of $H^{1}(A)$ (see [15, Ch. 6]).
-Only terms of total codimension two on $C\times A$ contribute to the first Chern class of $E$ after pushing forward.
-Rank of $E$
-The rank of $E$ comes from the degree-one term on $C$:
-rk(E)=d+r(1g).rk(E)=d+r(1-g).
-Since our slope assumption implies $H^{1}(C,V)=0$, Riemann–Roch gives
-$h^{0}(C,V)=d+r(1-g),$
-so we recover the rank formula obtained earlier in Proposition 3.3:
-$rk(E)=h^{0}(C,V).$
-First Chern class of $E$
-To compute $c_{1}(E)$, we isolate the codimension-two contribution coming from
-$r\cdot c_{1}(\mathcal{P})^{2}.$
-A classical computation (see [15, Ch. 6] and [6, Prop. 6.16]) shows that
-$p_{2*}\!\left(c_{1}(\mathcal{P})^{2}\right)=2\,\Theta.$
-Consequently,
-$c_{1}(E)=r\,\Theta.$
-Thus the determinant of $E$ is a positive multiple of the principal polarization.
-First Chern class of $F=E(\Theta)$
-Since twisting by $\Theta$ adds $rk(E)\Theta$ to the first Chern class, we obtain
-$c_{1}(F)=c_{1}(E)+rk(E)\,\Theta=\big(r+rk(E)\big)\Theta.$
-Chern classes of the kernel and the Ulrich bundle
-Recall the exact sequence constructed in Section 5:
-$0\longrightarrow M\longrightarrow G\longrightarrow F\longrightarrow 0,$
-where $G$ is a direct sum of degree-zero line bundles. Hence
-$c_{1}(G)=0,$
-and therefore
-$c_{1}(M)=-c_{1}(F).$
-Finally, the Ulrich bundle is defined by
-$U=M(m\Theta).$
-Twisting adds $rk(M)m\Theta$ to the first Chern class, giving
-$c_{1}(U)=c_{1}(M)+rk(M)m\Theta=\big(rk(M)m-r-rk(E)\big)\Theta.$
-Remark 8.1.
-In particular, all bundles arising in our construction have first Chern class proportional to the principal polarization $\Theta$.
-This reflects the intrinsic compatibility of the construction with the polarization of the Jacobian.
-It suggests that the resulting Ulrich bundles occupy a distinguished ray in the Néron–Severi space of $A$.
-9 Specialization to the Line Bundle
-Let $L$ be a line bundle on $C$. then $L$ be automatically semistable. Further we assume $\deg L>2g-2$. Thus we apply the above theory.
-Corollary 9.1.
-Let $L$ be a line bundle with $\deg L>2g-2$. Then $(J(C),m\Theta)$ carries an Ulrich bundle for all $m\geq g+1$.
-9.1 An upper bound for the Ulrich complexity
-In this subsection we compute rank of the Ulrich bundles produced by the construction of Section 5 in the case of line bundles.
-Let $L$ be a line bundle on $C$ of degree d>2g2d>2g-2, and set
-$E=\Phi_{\mathcal{P}}(a_{*}L),\qquad F=E(\Theta).$
-By Theorem 4.1, $F$ satisfies IT0. In particular, $F$ is continuously globally generated (CGG) in the sense of Pareschi–Popa (see [8, Remark 2.2(b)]).
-Hence there exist finitely many line bundles
-$\alpha_{1},\dots,\alpha_{N}\in\operatorname{Pic}^{0}(A)$
-such that the sum of the twisted evaluation maps yields a surjection
-$\varphi:G:=\bigoplus_{k=1}^{N}H^{0}(A,F\otimes\alpha_{k})\otimes\alpha_{k}^{-1}\longrightarrow F.$
-Let
-$0\longrightarrow M\longrightarrow G\stackrel{{\scriptstyle\varphi}}{{\longrightarrow}}F\longrightarrow 0$
-be the corresponding exact sequence, and define the Ulrich bundle
-$U:=M(m\Theta),\qquad m\geq g+1.$
-Computation of $h^{0}(A,F\otimes\alpha)$
-Since $F$ satisfies IT0, we have
-$H^{i}(A,F\otimes\alpha)=0\quad\text{for all }i>0\text{ and all }\alpha\in\operatorname{Pic}^{0}(A).$
-Therefore
-$h^{0}(A,F\otimes\alpha)=\chi(A,F\otimes\alpha)=\chi(A,F),$
-which is independent of $\alpha$.
-By Riemann–Roch on the abelian variety $A$ (using $\mathrm{td}(A)=1$),
-$\chi(A,F)=\frac{c_{1}(F)^{g}}{g!}.$
-From Section 8 we have
-$c_{1}(E)=\Theta,\qquad rk(E)=d+1-g.$
-Hence
-$c_{1}(F)=c_{1}(E)+rk(E)\Theta=(1+rk(E))\Theta.$
-Set
-n:=rk(E)=d+1g.n:=rk(E)=d+1-g.
-Then
-$c_{1}(F)=(1+n)\Theta.$
-Since $(\Theta^{g})=g!$ for a principally polarized abelian variety, we obtain
-$\chi(A,F)=\frac{(1+n)^{g}\Theta^{g}}{g!}=(1+n)^{g}.$
-Thus
-$h^{0}(A,F\otimes\alpha)=(1+n)^{g}\quad\text{for every }\alpha\in\operatorname{Pic}^{0}(A).$
-Rank of the Ulrich bundle
-By construction,
-$rk(G)=\sum_{k=1}^{N}h^{0}(A,F\otimes\alpha_{k})=N(1+n)^{g}.$
-Since $rk(F)=n$, it follows that
-$rk(U)=rk(M)=rk(G)-rk(F)=N(1+n)^{g}-n.$
-Substituting n=d+1gn=d+1-g, we obtain:
-Proposition 9.2.
-Let $L$ be a line bundle on $C$ of degree d>2g2d>2g-2. Then the Ulrich bundle $U$ constructed in Section 5 satisfies
-$rk(U)=N(d+2-g)^{g}-(d+1-g),$
-where $N$ is a positive integer depending on $F$, whose existence is guaranteed by continuous global generation. In particular,
-$\mathrm{uc}(J(C),m\Theta)\leq N(d+2-g)^{g}-(d+1-g).$
-Minimal degree choice
-If we choose the minimal possible degree d=2g1d=2g-1, then $n=g$ and
-$rk(U)=N(g+1)^{g}-g.$
-Then we get an upper bound
-$\mathrm{uc}(J(C),m\Theta)\leq N(g+1)^{g}-g.$
-Remark 9.3.
-The integer $N$ arises from the continuous global generation of $F$.
-While Pareschi–Popa guarantee the existence of such a finite $N$, no effective universal bound for $N$ is presently known in general.
-10 The induced moduli morphism
-In this section we discuss how the construction of Theorem A produces a morphism between moduli varieties. We retain the notation of the previous sections.
-10.1 Moduli spaces on the curve and on the Jacobian
-Fix integers $r\geq 2$ and $d$ with d>r(2g2)d>r(2g-2) and fix a line bundle $L_{0}\in Pic^{d}(C)$. Let
-$M_{C}:=M_{C}^{s}(r,L_{0})$
-denote the moduli variety of stable vector bundles on $C$ of rank $r$ and determinant $L_{0}$. It is a quasi-projective variety of dimension
-$\dim M_{C}=(r^{2}-1)(g-1)$
-(see [13, Ch. 4]).
-Fix $m\geq g+1$ and set $H=m\Theta$.
-Let $\mathbf{v}_{U}$ denote the numerical invariants (rank and Chern character) of the Ulrich bundles $U$ constructed in Section 5 from bundles $V$ of rank $r$ and degree $d$.
-Let
-$M_{A}:=M_{A}^{s}(\mathbf{v}_{U},H)$
-be the moduli variety of $H$-stable vector bundles on $A=J(C)$ with invariants $\mathbf{v}_{U}$.
-10.2 A universal family after passing to a cover
-Since $M_{C}$ need not carry a universal bundle, we pass to a generically finite cover.
-Lemma 10.1.
-There exists a quasi-projective variety $T$ and a generically finite surjective morphism
-$\pi:T\to M_{C}$
-together with a vector bundle $\mathcal{V}$ on $C\times T$ such that for each $t\in T$,
-$\mathcal{V}_{t}:=\mathcal{V}|_{C\times\{t\}}$
-represents the stable bundle corresponding to $\pi(t)$.
-Proof.
-This follows from the GIT construction of $M_{C}$ as a quotient of an open subset of a Quot scheme; see [13, Ch.
-4].
-Restricting to the stable locus and passing to a suitable cover yields a universal bundle on $C\times T$.
-∎
-10.3 Relative Fourier–Mukai transform
-Let $a:C\hookrightarrow A$ be the Abel–Jacobi embedding and $\mathcal{P}$ the normalized Poincaré bundle on $A\times\widehat{A}$, identified with $A$.
-Define
-$\mathcal{F}_{0}:=(a\times\mathrm{id}_{T})_{*}(\mathcal{V})\quad\text{on }A\times T,$
-and set
-$\mathcal{E}:=\Phi_{\mathcal{P}}(\mathcal{F}_{0})\quad\text{on }A\times T.$
-By the slope hypothesis $\mu(\mathcal{V}_{t})>2g-2$ and the arguments of Section 3, WIT0 holds fiberwise. After shrinking $T$ to a dense open subset, we may assume:
-  • (i)
-    $\mathcal{E}$ is a vector bundle on $A\times T$,
-  • (ii)
-    $\mathcal{F}:=\mathcal{E}(\Theta)$ satisfies IT0 fiberwise (Section 4),
-  • (iii)
-    $rk(\mathcal{E}_{t})$ is constant and equals $h^{0}(C,\mathcal{V}_{t})$.
-Openness of these properties follows from semicontinuity and cohomology and base change ([11, III.12], [8, §2]).
-10.4 Relative evaluation and kernel construction
-Fix the integer $N$ and general line bundles $\alpha_{1},\dots,\alpha_{N}\in Pic^{0}(A)$ as in Section 5.
-Define on $A\times T$:
-$\mathcal{H}_{k}:=p_{T*}(\mathcal{F}\otimes\alpha_{k}),\qquad\mathcal{G}:=\bigoplus_{k=1}^{N}p_{T}^{*}\mathcal{H}_{k}\otimes\alpha_{k}^{-1},$
-where $p_{T}:A\times T\to T$.
-Since $\mathcal{F}$ satisfies IT0, each $\mathcal{H}_{k}$ is locally free and base change holds.
-The fiberwise evaluation maps
-$G(V_{t})\twoheadrightarrow F(V_{t})$
-constructed in Section 5 assemble into a morphism
-$\mathrm{ev}:\ \mathcal{G}\longrightarrow\mathcal{F}$
-which is surjective on a dense open subset of $T$ by continuous global generation ([8, Remark 2.2]).
-After shrinking $T$, assume $\mathrm{ev}$ is surjective. Define the kernel
-$0\longrightarrow\mathcal{M}\longrightarrow\mathcal{G}\longrightarrow\mathcal{F}\longrightarrow 0$
-and set
-$\mathcal{U}:=\mathcal{M}(m\Theta).$
-Then $\mathcal{U}$ is flat over $T$ and its fiber over $t\in T$ is precisely the Ulrich bundle $U(V_{t})$ constructed in Section 5.
-10.5 Stability in families
-By Section 7, for general evaluation data the bundle $U(V_{t})$ is $\mu_{H}$-stable. Since stability is open in flat families ([13, Thm. 2.3.1]), there exists a dense open subset
-$T^{\mathrm{st}}\subset T$
-such that for all $t\in T^{\mathrm{st}}$, $\mathcal{U}_{t}$ is $H$-stable.
-10.6 A morphism of moduli varieties
-By the universal property of the coarse moduli space of stable bundles, the family $\mathcal{U}$ over $T^{\mathrm{st}}$ induces a morphism
-$\Phi:\ T^{\mathrm{st}}\longrightarrow M_{A}^{s}(\mathbf{v}_{U},H),\qquad t\longmapsto[\mathcal{U}_{t}].$
-This is a morphism of quasi-projective varieties.
-10.7 Geometric consequences
-Retaining all the notations in this section we state the geometric consequences as a theorem.
-Theorem 10.2.
-Fix integers $r\geq 2$ and d>r(2g2)d>r(2g-2), and let $M_{C}$ denote the moduli space of stable vector bundles on $C$ of rank $r$ and degree $d$.
-For $m\geq g+1$, let $M_{A}$ be the moduli space of $H$-stable bundles on $(A,m\Theta)$ with the numerical invariants of the Ulrich bundles constructed in Theorem A.
-Then there exists a quasi-projective variety $T$, a generically finite surjective morphism
-$\pi:T\longrightarrow M_{C},$
-and a dense open subset $T^{\mathrm{st}}\subset T$ such that the Ulrich construction is compatible with families and induces a morphism
-$\Phi:T^{\mathrm{st}}\longrightarrow M_{A}.$
-In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$.
-Proof.
-By Lemma 10.1, there exists a quasi-projective variety $T$ and a generically finite surjective morphism
-$\pi:T\longrightarrow M_{C},$
-together with a vector bundle $\mathcal{V}$ on $C\times T$ whose fiber $\mathcal{V}_{t}$ represents the point $\pi(t)\in M_{C}$.
-Consider the family $(a\times id_{T})_{*}\mathcal{V}$ on $A\times T$ and apply the relative Fourier–Mukai transform with kernel the Poincaré bundle.
-The cohomology-and-base-change arguments of Sections 3 and 4, applied fiberwise using the hypothesis d>r(2g2)d>r(2g-2), show that after possibly shrinking $T$, the transform
-$\mathcal{E}:=\Phi_{\mathcal{P}}\big((a\times id_{T})_{*}\mathcal{V}\big)$
-is locally free on $A\times T$ and commutes with base change. Moreover, the twist $\mathcal{F}:=\mathcal{E}(\Theta)$ satisfies the $\mathrm{IT}_{0}$ property fiberwise (Section 4).
-Since $\mathcal{F}$ is $\mathrm{IT}_{0}$ on each fiber, it is continuously globally generated in the sense of Pareschi–Popa.
-Fix line bundles $\alpha_{1},\dots,\alpha_{N}\in Pic^{0}(A)$ as in Section 5 and form the relative evaluation morphism
-$\mathrm{ev}:\ \mathcal{G}:=\bigoplus_{k=1}^{N}p_{T}^{*}\!\big(p_{T*}(\mathcal{F}\otimes\alpha_{k})\big)\otimes\alpha_{k}^{-1}\longrightarrow\mathcal{F},$
-where $p_{T}:A\times T\to T$ is the projection. By construction and upper semicontinuity of cohomology, after shrinking $T$ further if necessary, the morphism $\mathrm{ev}$ is surjective.
-Let $\mathcal{M}:=\ker(\mathrm{ev})$ and define
-$\mathcal{U}:=\mathcal{M}(m\Theta)\quad\text{on }A\times T.$
-Then $\mathcal{U}$ is flat over $T$, and for every $t\in T$, the fiber $\mathcal{U}_{t}$ coincides with the Ulrich bundle on $(A,m\Theta)$ constructed from $\mathcal{V}_{t}$ in Theorem A (Section 5).
-In particular, each $\mathcal{U}_{t}$ is Ulrich with respect to $H=m\Theta$.
-Let $T^{\mathrm{st}}\subset T$ be the locus where $\mathcal{U}_{t}$ is $H$-stable.
-Since slope-stability is an open property in flat families of coherent sheaves (see [13, Thm.
-2.3.1]), the subset $T^{\mathrm{st}}$ is Zariski open.
-By the stability results of Section 8 (for general evaluation data and sufficiently positive $m$), this locus is nonempty; hence $T^{\mathrm{st}}$ is dense in $T$.
-By the universal property of the coarse moduli space $M_{A}=M_{A}^{s}(v_{U},H)$, the family $\mathcal{U}|_{A\times T^{\mathrm{st}}}$ induces a morphism
-$\Phi:T^{\mathrm{st}}\longrightarrow M_{A},\qquad t\longmapsto[\mathcal{U}_{t}],$
-see [13, Cor. 4.6.6]. This proves the existence of $\pi$, $T^{\mathrm{st}}$, and $\Phi$ as stated.
-Finally, since $\pi:T\to M_{C}$ is generically finite and dominant, one has
-$\dim T=\dim M_{C}=(r^{2}-1)(g-1),$
-and the same holds for the dense open subset $T^{\mathrm{st}}$.
-Therefore the image $\Phi(T^{\mathrm{st}})\subset M_{A}$ is swept out by a family of stable Ulrich bundles parametrized by a variety of dimension $(r^{2}-1)(g-1)$.
-In particular, $M_{A}$ contains positive-dimensional families of stable Ulrich bundles arising from stable bundles on $C$.
-∎
-Acknowledgements
-I thank IISER Berhampur for its cohesive research environment and sustained support. I sincerely thank in advance to all readers of this arXiv version for kind comments/remarks/suggestions (if any).